A Complete Bibliography of The American Statistician: 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

07 October 2024 Version 1.01

Title word cross-reference

χ^{2} [Nat59]. t [Ban57, Cur57, Dos57, Kna57]. t^{2} [Dwy58].

110th [Ano50a]. 111th [Ano51a]. 112th [Ano52a]. 113th [Ano53a]. 114th [Ano54a]. 115th [Ano55x]. 116th [Ano56w]. 117th [Ano57u]. 118th [Ano58x]. 119th [Ano59-28]. 125th [Mac59]. 16th [Ano53b]. 1950 [Eck53]. 1951 [Ano51j, Dor51]. 1952 [Hoa52]. 1954 [Ano54b, CG58]. 1955 [Dal55b]. 1959 [Ric59].

701 [Ano56g, BQ56]. 704 [Ano56g, BQ56]. 75th [Qua51].

Abraham [Dal55b, Ano51-31, Hot51]. Abstracts [Ano50b, Ano50-27]. Accepted [Ric52]. Accident [Ger51]. Activities [Alp53, Ano50c, Ano51c, Ano58m, Ano59m, Ano59n, Ano59o, Ano59p, Ano59q, BGS^{+}55, BDG57, BTG^{+}58, BFR59, BD59, BWW^{+}52, DKC^{+}56, Gal52, GGL^{+}56, GSP^{+}58, HGG^{+}53, HZR^{+}52, JWLW53, LRP^{+}57, LDL^{+}54, LGD^{+}55, LGLH55, LYG^{+}56, Leo52, Lin52, LBW^{+}57, MDY^{+}57, MYG^{+}53, MLRB54, MSLC57, MS59, MM59, Mil52, Mor50a, MKB^{+}54, MSBH59, RLV^{+}55, RJKP58, SCJ59, SMW^{+}54, SK58, SLCM59, SLK^{+}58, SWHS53, SDF^{+}59, SLW^{+}56, SPE54, SWL56, YLMM53, YHY^{+}55]. Actuarial [Ano51i, FP51b]. Ad [Ano56x, CCF^{+}54]. Administration [Ano52o]. Admissibility [Fer56b]. Advanced [Mah50]. Advisory [Ano52z, CDK^{+}53, CCD^{+}53, Ano54-27]. Aggregate [Fer56a]. Agricultural [Vol53]. Aid [WD50]. Amendment [Wei54]. American [Ano50a, Ano50b, Ano50-29, Ano51a, Ano52a, Ano52z, Ano53a, Ano54a, Ano55x, Ano56w, Ano57u, Ano58x, Ano59-28, Ano52i, Ano54b, CCFM53, Fit55, Fit56, Ger51, Joh53, RB59, WR55]. among [Mah58]. Analysis [Ano56g, BQ56, Kin52]. Analyzing [Moo54]. Anniversary [Mac59, Qua51]. Annual [Ano50a, Ano51a, Ano52a, Ano53a, Ano53b, Ano54a, Ano55x, Ano56w, Ano57u, Ano58x, Ano59-28, Dor51, Mor50a, Sho59]. Answer [Don51, Spr59]. Answers [CSDP52, CSP^{+}53b, CSP^{+}53a, CSC^{+}53, Dun55, Edw59, MRH50, MM50, Mos50b, Mos50c, MMS51, OCSP52, Rub54a, Rub54b, Rub54c, RE55, RJ55, RS55, RR55, Rub56a, Rub56b, RS56, Rub56c, Rub56d, RJ57a, Rub57a, Rub57b, RJ57b, RD57, RH58, Rub58a, Rub58b, RS58, Rub58c, RSPB59, Rub59, RH59b, RH59a, RH59c]. Apparatus [Sch53]. Applications [GL54]. Applied [Day50]. Appraisal [Ber53]. Areas [GBF^{+}54, WBB^{+}56]. Arranged [Ano54i]. Arts [FM56a]. ASA [CDK^{+}53, CCD^{+}53, GBF^{+}54, MWS54, SWW53]. Aspects [FM56b, Lof52]. Association [Ano50c, Ano51c, Ano52z, Ano54b, Ano50a, Ano50i, Ano51a, Ano51h, Ano52a, Ano52h, Ano52i, Ano52x, Ano53a, Ano53h, Ano54a, Ano54h, Ano55x, Ano56w, Ano57u, Ano58x, Ano59-28, CCFM53, Ger51, RB59, WR55]. Automatic [Spr56]. Availability [Ano59-29]. Average [Ano52b]. Avoiding [Com50].

Back [Ano55a, Ano55b, Ano55c, Ano56a, Ano58a, Ano58b, Ano59a, Ano59b]. Balance [Pol59]. Base [Ric51]. Basic [Dwy57]. Become [Yat50]. Below [DK50, Bat50]. Benedict [Hoa52]. Berkeley [Ano54b]. Bias [KH59]. Binary [Har53]. Blood [MDM51]. Both [Har58]. Brains [Ber50]. Branch [Lin52]. Budget [Ano52z, Ano54-27, CDK^{+}53, CCD^{+}53]. Bureau [Ano52z, Mil52, Moo50, Ano54-27, CDK^{+}53, CCD^{+}53, Dal57, Mos50a]. Business [Hoa51, Moo54, WW50]. By-Laws [Wei54].

Calendar [Arm50a, Arm50b, Arm51]. California [Ano54b]. Canada [MS59, MM59, SCJ59]. Canadian [Car54, Leo54, Mar54]. Cancer [Dor54]. Card [Har53]. Career [Ano52-27]. Celebrates [Qua51]. Census [CG58, GBF^{+}54, Han52a, Han52b, Lon52a, Lon52b, Old57, Ull59, WBB^{+}56, Yat50, Dal57, Eck53]. Central [Ano52o]. Century [Fit57]. Change [Don51]. Changes [EBL52, Spu52]. Chapel [Ano50p]. Chapter [Ano50d, Ano50e, Ano50f, Ano50g, Ano50h, Ano51d, Ano51e, Ano51f, Ano51g, Ano52c, Ano52d, Ano52e, Ano52f, Ano52g, Ano53c, Ano53d, Ano53e, Ano53f, Ano53g, Ano54c, Ano54d, Ano54e, Ano54f, Ano54g, Ano55d, Ano55e, Ano55f, Ano55g, Ano55h, Ano56b, Ano56c, Ano56d, Ano56e, Ano56f, Ano57a, Ano57b, Ano57c, Ano57d, Ano57e, Ano58c, Ano58d, Ano58e, Ano58f, Ano58g, Ano59c, Ano59d, Ano59e, Ano59f, Ano59g, FP51a]. Character [Eck53]. Characteristics [RB59]. Chart [Hae50a, Hae50b, Stu50]. Chicago [FP51a, Sho59, Wal51b]. Chile [Rot58]. China [Kle59]. Classification [Bur59]. Clinical [Gre59, MDM51]. Co [Car54, Leo54, Mar54]. Co-Ordination [Car54, Leo54, Mar54]. Coding [Har53]. Coin [Spr56]. Collecting [Alp53]. College [Bat50, DK50]. Colleges [Ban58, Cha55a, Fit55, Har58, Joh53]. Color [Hae50a]. Column [Hoa58a, Hoa58b, Hoa58c, Lik59a, Lik59b, Lik59c, Lik59d, Lik59e]. Combined [Sag53]. Comment [Moo55, Moo57]. Comments [Car54, Cre52, Leo54]. Commercial [Sim56]. Committee [Ano52z, Ano54-27, Ano56x, CDK^{+}53, CCD^{+}53, CCF^{+}54, GBF^{+}54, ADG^{+}52, Ano50-28, Ano51-29, Ano51-30, Ano51b, Ano53y, CCD^{+}59, CLE52, CDK^{+}51, CCFM53, Dor51, Fre53, Ger51, Kel53, WRW54, WBB^{+}56]. Committees [Ano50i, Ano51h, Ano52h, Ano53h, Ano54h, Ano51-29]. Communist [Kle59]. Company [Kin52, Wri52, Bag52]. Computation [Nat59]. Computational [Dwy58]. Computations [Col59, Dal57, Dwy57, Fer56b, Shi55]. Computers [CG58]. Conceptions [Ric52]. Concerning [Bat50]. Conduct [Gre59]. Conference [Ano52o, Ano50o, Sho59]. Congress [Ano50b, Ano50-29]. Conscience [Bro52, Fre52]. Constitution [Ano52i, EBL52]. Constructing [Edw59]. Consumer [Dau54, Jon54, Jaf54]. Content [Ull59]. Control [Ano50-31, Koy51, Lit50, Lit51, Wix52]. Cooperation [Mos50a]. Cooperative [Ano54i, Gre59]. Coordination [Alp53, Ric50]. Correction [Ano56g]. Correlation [Ano56g, BQ56, Edw59, WD50]. Correlations [Mah58]. Council [MWS54, SWW53]. Course [Mah50]. Courses [Ano50-31, Ano51i, Fit56, Ros50, Spu52]. Court [Fer56b]. Credit [Dau54, Jon54]. Critical [Gol53]. Cultural [Ano50-27]. Current [Jaf54, KDOR51, MGCK51, MP51, OCMH50, RSS^{+}52]. Curricula [Spu52]. Curriculum [Ano51i, FM56a]. Cycles [Moo54].

Daily [Sma54]. Data [Alp53, Ano56g, Arm50a, Arm50b, BQ56]. December [Ano54b]. Defense [FM55]. Degree [Har51]. Delivered [Mac59]. Demographic [Ano51i, FP51b, Lin52]. Demonstrations [Spr56]. Denial [KS58]. Description [Vin52]. Design [Day54, Gol53]. Designs [Ken59]. Desirable [Fra52, Gul52, Tep52]. Development [Cre52, Der52, Lit50, Lit51]. Developments [Joh53]. Device [BL51]. Diffusion [Bro55, Moo55, Moo57, Val57a, Val57b]. Dimension [Hae51]. Dinner [Mac59]. Directors [Ano52z]. Discussion [Dal55a, Fra52, Fre52, Han52b, Kis52, Lon52a, Mor52, Tep55]. Distribution [Abr51, CPT56, Cox52]. Division [Ano52z]. Dr. [Kis52]. Draft [Ano50-27].

Early [Fit55, Fit56]. Econometric [Ano51-28, Ano52y]. Economic [Bol56, Hoa51, Lof52, Mil52, Moo50, Ste55b, Tep55, WW50]. Economics [Gla51, Vin52]. Economists [Ano52b]. Editor [Ano51p, DC54, HS51, KKP^{+}54, Net55, Pet50, Sie52, Smi56, Sta59]. Educational [Ano50-27, But59, Gul52]. Effective [Bon50]. Effects [Arm50a, Arm50b]. Egypt [Sar57]. Eight [Ber50]. Election [Ano54j, Wei54]. Elections [Ano50-28, Ano51-30, Ano53y, CLE52, WRW54]. Electronic [Ano56g, BQ56, CG58, Har53]. Emanuel [Tho53]. Empirical [CPT56]. Employees [Gul52]. Employment [Ano54i, Dal55a, Sim56]. Engineering [Ano50-31, Day54]. Engineers [Bur50, Olm51]. Enumeration [GBF^{+}54, WBB^{+}56]. Errors [Eck53, Man51]. Especially [Mor50a]. Estimate [CPT56]. Estimating [Sma54]. Estimation [Man51]. Evidence [Dem58]. Examined [Ric52]. Experience [CG58]. Experiment [Ken59, Qua51]. Experiments [Day54]. Experts [Yat50]. Explore [Ano56x]. Extent [Eck53]. Extreme [GL54]. Extreme-Value [GL54]. Ezra [Vin52].

Faculty [Ano51i, FP51b]. Federal [Ano59-29, BGS^{+}55, BDG57, Bow57, BTG^{+}58, BFR59, BD59, BWW^{+}52, DKC^{+}56, Gal52, GGL^{+}56, GSP^{+}58, HGG^{+}53, HZR^{+}52, JWLW53, LRP^{+}57, LDL^{+}54, LGD^{+}55, LGLH55, LYG^{+}56, LBW^{+}57, MDY^{+}57, MYG^{+}53, MLRB54, MSLC57, MKB^{+}54, MSBH59, Ric50, RLV^{+}55, RJKP58, SMW^{+}54, SK58, SLCM59, SLK^{+}58, SWHS53, SDF^{+}59, SLW^{+}56, SPE54, SWL56, YLMM53, YHY^{+}55, Alp53, Fut58, MGCK51, MP51, RSS^{+}52, Sag53, Wic54]. Fellows [Ano54j, CDK^{+}51, Kel53, Wei54]. Field [Gre59]. Fielding [Ric59]. Fields [Har53]. Films [But59]. Financial [WR55, Wix52]. Fit [Rif59]. Flipper [Spr56]. Forces [Ric53b]. Forecasting [FP51a]. Forecasts [Ano52b, Kin52]. Formula [Ban57]. Formulas [Dwy58]. Frantisek [Ano51j]. Frequencies [FM55]. Front [Ano50j, Ano50k, Ano50l, Ano50m, Ano50n, Ano51k, Ano51l, Ano51m, Ano51n, Ano51o, Ano52j, Ano52k, Ano52l, Ano52m, Ano52n, Ano53i, Ano53j, Ano53k, Ano53l, Ano53m, Ano54k, Ano54l, Ano54m, Ano54n, Ano54o, Ano55i, Ano55j, Ano55k, Ano55l, Ano55m, Ano56h, Ano56i, Ano56j, Ano56k, Ano56l, Ano57f, Ano57g, Ano57h, Ano57i, Ano57j, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l]. Further [Cur57, Dos57]. Fuzes [Day50].

General [Ban57, Ric51]. General-Purpose [Ric51]. Generally [Ric52]. German [Hug55]. Girshick [Dal55b]. Glaser [Vin52]. Gleichschaltung [Hug55]. Glenn [Har50]. Going [Ano50-30]. Goldenweiser [Tho53]. Government [Alp53, Fut58, KDOR51, MS59, MM59, MGCK51, MP51, OCMH50, RSS^{+}52, SCJ59]. Graphic [Rif59]. Group [Cur57].

Helps [FP51a]. Hiawatha [Ken59]. High [Ros50]. Hill [Ano50p]. Hoc [Ano56x, CCF^{+}54]. Honor [Ano51-31]. Honoring [Ano54-28]. Horizontal [Har53]. Hundred [Ber50]. Hypotheses [Spr59].

IBM [Ano56g, BQ56]. III [CPT56, Com50]. Illinois [Ano50o]. Implications [MDM51]. Improvement [Bow57]. Income [Abr51, Cre52, Der52, Ger53]. Index [Ano52b, Moo57, Val57a, Val57b]. Indexes [Bro55, Moo55, Ric51]. Indicators [Mah58]. Industrial [Ano52b, Ger53, Sim56]. Industry [Har58]. Influence [Stu50]. Influences [Arm51]. Information [Gil58]. Institute [Ano53b, Ano50p]. Institutional [Ano51b]. Instruction [Cha55a]. Instructions [Dwy57]. Instructors [Ano51i]. Intend [Yat50]. Inter [Ano50b, Ano50-29]. Inter-American [Ano50b, Ano50-29]. Interested [Ber50]. Interindustry [Gla51, Vin52]. International [Ano58m, Ano59m, Ano59n, Ano59o, Ano59p, Ano59q, Lof52, Mos50a]. Interpretation [Fer56a]. Intra [Wri52]. Intra-Company [Wri52]. Introduction [Ric53a]. Introductory [FM56a]. IV [Hay50].

J. [Ano52b]. Japan [Koy51]. Jersey [Wri52]. John [Har50]. Joint [Stu50].

Kelly [Day53].

Labor [Mos50a]. Lantern [Bon50]. Large [Dal57]. Laws [KS58, Wei54]. Lead [Moo57, Val57a, Val57b]. Leading [Ano52b]. Least [Rif59]. Legal [Dem58]. Legislation [Ric50]. Letter [Ano53x, Net55, Pet50, Smi56, Sta59]. Letters [Ano51p, DC54, HS51, KKP^{+}54, Sie52, SK51]. Level [Bat50, DK50]. Levels [Gul52]. Liberal [FM56a]. Life [Bag52]. Limitations [Dau54]. Limits [Woy54]. Linder [Kis52]. Literacy [Wal51a]. Livingston [Ano52b]. Loan [Bag52]. Lung [Dor54].

M [Har50]. Machine [Har53]. Machines [Ano56g, BQ56]. made [Ano52b]. Mail [Sma54]. Manufactures [CG58]. Map [Com50]. Marketing [Ald53]. Materials [Ano59-29]. Math [Mur56]. Mathematical [Ano53b]. Mathematics [Gul52, Woy54]. Matter [Ano50j, Ano50k, Ano50l, Ano50m, Ano50n, Ano51k, Ano51l, Ano51m, Ano51n, Ano51o, Ano52j, Ano52k, Ano52l, Ano52m, Ano52n, Ano53i, Ano53j, Ano53k, Ano53l, Ano53m, Ano54k, Ano54l, Ano54m, Ano54n, Ano54o, Ano55a, Ano55b, Ano55c, Ano55i, Ano55j, Ano55k, Ano55l, Ano55m, Ano56a, Ano56h, Ano56i, Ano56j, Ano56k, Ano56l, Ano57f, Ano57g, Ano57h, Ano57i, Ano57j, Ano58a, Ano58b, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano59a, Ano59b, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l]. Meaning [Dal55a]. Means [Man51, Stu50]. Measurement [Sim56, Ste55a]. Mechanical [Ber50]. Meeting [Ano50a, Ano51a, Ano52a, Ano53a, Ano53b, Ano54a, Ano54b, Ano55x, Ano56w, Ano57u, Ano58x, Ano59-28, Dor51, Mor50a]. Meetings [Sie50]. Members [Ano50v, Ano50w, Ano50x, Ano50y, Ano50z, Ano51w, Ano51x, Ano51y, Ano51z, Ano51-27, Ano52v, Ano52w, Ano53s, Ano53t, Ano53u, Ano53v, Ano53w, Ano54v, Ano54w, Ano54x, Ano54y, Ano54z, Ano55s, Ano55t, Ano55u, Ano55v, Ano55w, Ano56r, Ano56s, Ano56t, Ano56u, Ano56v, Ano57p, Ano57q, Ano57r, Ano57s, Ano57t, Ano58s, Ano58t, Ano58u, Ano58v, Ano58w, Ano59w, Ano59x, Ano59y, Ano59z, Ano59-27]. Membership [Ano53x, RB59, WR55]. Memberships [Ano51b]. Method [Hot58, Rif59]. Methodology [Ano52o]. Methods [Day50, GL54, Yat50]. Meyer [Dal55b]. Mid [Joh53]. Mid-American [Joh53]. Midwest [Sho59]. Missouri [Har51]. Mobility [Gol53]. Model [WD50]. Monthly [Arm50a, Arm50b]. Moore [Mah58]. Mortgage [Bag52]. Mr. [Vin52]. Multiple [Ano56g, BQ56]. Must [Moo57, Val57a, Val57b].

National [Cre52, Der52, Ger53, Dor50, Mil52, Moo50]. Nations [Ano50-27, Leo52, Lin52]. Nature [Lor53]. Naval [Day54]. News [Ano50v, Ano50w, Ano50x, Ano50y, Ano50z, Ano50q, Ano50r, Ano50s, Ano50t, Ano50u, Ano51q, Ano51w, Ano51x, Ano51y, Ano51z, Ano51-27, Ano51r, Ano51s, Ano51t, Ano51u, Ano51v, Ano52p, Ano52v, Ano52w, Ano52q, Ano52r, Ano52s, Ano52t, Ano52u, Ano53s, Ano53t, Ano53u, Ano53v, Ano53w, Ano53n, Ano53o, Ano53p, Ano53q, Ano53r, Ano54v, Ano54w, Ano54x, Ano54y, Ano54z, Ano54q, Ano54r, Ano54s, Ano54t, Ano54u, Ano55s, Ano55t, Ano55u, Ano55v, Ano55w, Ano55n, Ano55o, Ano55p, Ano55q, Ano55r, Ano56r, Ano56s, Ano56t, Ano56u, Ano56v, Ano56m, Ano56n, Ano56o, Ano56p, Ano56q, Ano57p, Ano57q, Ano57r, Ano57s, Ano57t, Ano57k, Ano57l, Ano57m, Ano57n, Ano57o, Ano58s, Ano58t, Ano58u, Ano58v, Ano58w, Ano58n, Ano58o, Ano58p, Ano58q]. News [Ano58r, Ano59w, Ano59x, Ano59y, Ano59z, Ano59-27, Ano59r, Ano59s, Ano59t, Ano59u, Ano59v, Net52]. Nineteenth [Fit57]. No [Don51]. Nongovernmental [Ano59-29]. Nonparametric [Sie57]. Nonresponse [KH59]. Normal [MDM51]. Note [Cur57, Dos57, Kle59, Kna57, Nat59, Rif59, Ros50]. Notes [Ano50d, Ano50e, Ano50f, Ano50g, Ano50h, Ano51d, Ano51e, Ano51f, Ano51g, Ano52c, Ano52d, Ano52e, Ano52f, Ano52g, Ano53c, Ano53d, Ano53e, Ano53f, Ano53g, Ano54c, Ano54d, Ano54e, Ano54f, Ano54g, Ano55d, Ano55e, Ano55f, Ano55g, Ano55h, Ano56b, Ano56c, Ano56d, Ano56e, Ano56f, Ano57a, Ano57b, Ano57c, Ano57d, Ano57e, Ano58c, Ano58d, Ano58e, Ano58f, Ano58g, Ano59c, Ano59d, Ano59e, Ano59f, Ano59g].

Observation [Hol58]. Obtaining [Rif59]. Occupations [Gol53]. Office [Ano54-27, Lin52]. Ogburn [Ric59]. Oil [Wri52]. Oldest [Qua51]. Once [Pol59]. One [Don51, Jaf54, Bag52]. Open [Ano53x]. Operation [Sag53, Sch53]. Operations [Lat54]. Opinion [Ano56x, Bat50]. Order [Sma54]. Ordination [Car54, Leo54, Mar54]. Organisms [KS58]. Organization [Ano52z, Ano54-27, Gul52, Hol58, Sch53, Ano50-27]. Organizational [Dal57]. Outline [Gil58].

Panel [Cox52]. Paper [Ano50-30, Kis52]. Partial [WD50]. Pearson [CPT56]. People [Ber50]. Period [Ric51]. Permanent [Ano54-27]. Philippine [Ano52x]. Philosophy [FM56b, Hot58]. Placement [Ano54i, Dor50]. Planning [Wix52]. Plans [Sie50]. Point [Fut58, Hay50]. Policy [Ano52z, Ano54-27, Bol56, CCD^{+}59, Ste55b, Tep55, CDK^{+}53, CCD^{+}53]. Population [Jaf54, Lor53, Ull59]. Practical [Stu50]. Predetermined [Edw59]. Preparation [Bon50, Sta50]. Present [Ano50-30]. Presentation [Com50, Hae50a]. Presenting [Stu50]. President [Hoa58a, Hoa58b, Hoa58c, Lik59a, Lik59b, Lik59c, Lik59d, Lik59e]. Pressure [MDM51]. Probability [Dem58]. Problem [Com50, Don51, FM56b]. Problems [Bur59, Car54, Cha55b, Dal57, Edw59, Leo54, Mar54]. Procedure [KH59]. Processing [Ano56g, BQ56, CG58]. Production [Ano52b, Sag53, Bow52]. Professional [Mah50, Yat50]. Professor [Gla52]. Program [Ano50-27, Ano51-28, Ano52y, Cha55b, Dor51, Fra52, Tep52, WW50, Ano55x, Ano56w, Ano57u, Ano58x, Ano59-28, Hay50, Mos50a]. Programs [Joh53]. Projectiles [Day50]. Projects [KDOR51, MGCK51, MP51, OCMH50, RSS^{+}52]. Proposal [Dor50]. Proposals [Wic54]. Proposed [Wei54]. Providing [Cha55b]. Proximity [Day50]. Pseudo [Hae51]. Publications [CCD^{+}59]. Purdue [Bur50]. Purpose [Ric51].

Quality [Ano50-31, Koy51, Lit50, Lit51]. Quantitative [Ros50]. Questionnaire [CCD^{+}59]. Questions [CSDP52, CSP^{+}53b, CSP^{+}53a, CSC^{+}53, Dun55, MRH50, MM50, Mos50b, Mos50c, MMS51, OCSP52, Rub54a, Rub54b, Rub54c, RE55, RJ55, RS55, RR55, Rub56a, Rub56b, RS56, Rub56c, Rub56d, RJ57a, Rub57a, Rub57b, RJ57b, RD57, RH58, Rub58a, Rub58b, RS58, Rub58c, RSPB59, Rub59, RH59b, RH59a, RH59c, Ste55a].

Random [Pol59]. Ranking [Hae50b]. Rapid [Man51]. Ratio [Fer56a]. Reading [Vin52]. Receipts [Sma54]. Recommendations [Ano59-29]. Reducing [KH59]. Reflections [Gla52, Vin52]. Regional [Sie50, Ano54b]. Regression [Ano56g, BQ56]. Rejoinder [Lon52b, Val57b]. Related [Dal57]. Relating [Ano50-27]. Relationship [Dor54]. Reliability [Ber53, Dal55a, Ger53, Vol53]. Replacement [KH59]. Report [ADG^{+}52, Ano50-28, Ano51-29, Ano51-30, Ano51b, Ano52z, Ano53y, Ano56x, Bow52, CCD^{+}59, CDK^{+}53, CCD^{+}53, CCF^{+}54, CLE52, CDK^{+}51, CCFM53, Cox52, Dor51, Fre53, Ger51, GBF^{+}54, Kel53, Mil52, MWS54, Sie50, SWW53, WRW54, WBB^{+}56, Wix52]. Reports [Sta50]. Representative [Ger51]. Representatives [Ano50i, Ano51h, Ano52h, Ano53h, Ano54h, MWS54, SWW53]. Research [Ald53, Ano59-29, Gla51, Gol53, Gul52, Lat54, MWS54, SWW53, Vin52, Mil52, Moo50]. Reservations [Gla52]. Resolution [Ano51-31, Ano54-28]. Resolutions [Ano50b]. Results [CCD^{+}59]. Review [Alp53]. Revision [Ric51]. Revisions [Jon54]. Rice [Ano54-28]. Role [Bol56, Wei50]. Rome [Ano51i, FP51b]. Royal [Mac59]. Russia [Vol53].

Sales [Kin52]. Samples [Man51]. Sampling [BL51, Bur59, Dem58, Sta50, Spr56, Yat50, Mah50]. Samuel [K.55]. Saurino [Hoa52]. Savings [Fer56a]. Scale [Dal57]. Scatter [Stu50]. Scheduling [Bow52]. School [Ros50]. Science [Hot58, Lat54, MWS54, SWW53]. Sciences [Ano51i, FP51b, Wal51a, Ano50-27]. Scientific [Ano50-27, FM56b, Ric52]. Seasonal [Arm51, Fer56b, Shi55]. Second [Ano50b, Ano50-29]. Section [Ano54p, Hoa51, Mor50b, Bow52, Wix52, WW50]. Series [Arm50a, Arm50b, Mah58]. Service [Ano54i, Dor50]. Shaping [Bol56]. Sheet [Spr59]. Shifts [Arm50a, Arm50b]. Simple [Edw59, Stu50]. Simplified [Col59, Hae50b]. Skewness [CPT56]. Sliderule [Day53]. Slides [Bon50]. Small [Com50, Man51]. Social [ADG^{+}52, Ano50-27, Ano54p, Lin52, Mor50a, Old57, Ric53b, Wal51a, MWS54, SWW53]. Societies [Fit57]. Society [Ano51-28, Ano52y, Mac59]. Some [CG58, Cre52, FM56b, GL54, Ste55a]. South [Cha55a]. Soviet [Ber53, Ger53, Lor53, Ric52, Sch53, Vol53]. Speech [Mac59]. Sponsored [Alp53]. Squares [Rif59]. Standard [Man51, Wri52]. Standards [Ano52z, Cou52, Dem58, Ger51, Han52a, Han52b, Lon52a, Lon52b, Mor52, Ano54-27, Ano56x, CCF^{+}54]. Standpoint [Ban58]. State [Com50, Sag53]. States [Fit57, Lit50, Lit51, Sim56, Wic54]. Station [Qua51]. Statistical [Ald53, Ano50a, Ano50b, Ano50-29, Ano51a, Ano51i, Ano52a, Ano52i, Ano52o, Ano52x, Ano52z, Ano53a, Ano54a, Ano54b, Ano54-27, Ano55x, Ano56w, Ano57u, Ano58m, Ano58x, Ano59m, Ano59n, Ano59o, Ano59p, Ano59q, Ano59-28, Ano59-29, Bat50, BGS^{+}55, BDG57, BTG^{+}58, BFR59, BD59, BWW^{+}52, CDK^{+}53, CCD^{+}53, Car54, CCF^{+}54, Cou52, CCFM53, Dal57, Day50, DKC^{+}56, DK50, Dwy57, FP51b, Fit57, Fit56, FM56b, Gal52, Ger51, GGL^{+}56, GSP^{+}58, Han52a, Han52b, Har53, HGG^{+}53, Hol58, HZR^{+}52, Hot58, Hug55, Joh53, JWLW53, KDOR51, Kle59, Koy51, LRP^{+}57, LDL^{+}54, LGD^{+}55, LGLH55, LYG^{+}56, Leo52, Leo54, Lin52, LBW^{+}57, Lit50, Lit51, Lon52a, Lon52b, Mac59, Mah50, Mah58, MDY^{+}57, Mar54, MYG^{+}53, MLRB54, MSLC57, MS59, MM59, MGCK51, Mor52, MKB^{+}54, MSBH59, MP51]. Statistical [Olm51, OCMH50, Ric50, Ric52, RLV^{+}55, RSS^{+}52, RJKP58, RB59, SCJ59, Sch53, SMW^{+}54, SK58, SLCM59, SLK^{+}58, SWHS53, SDF^{+}59, SLW^{+}56, SPE54, SWL56, Wal51a, WR55, Wic54, YLMM53, YHY^{+}55]. Statistician [Bro52, Fre52, K.55, Wei50]. Statisticians [Mor50b, Nic54, Old57, Yat50]. Statistics [Ano50p, Ano50o, Ano50-31, Ano52-27, Ano54p, Bag52, Ber53, Bol56, Bow57, Bur50, But59, Cha55a, Cha55b, Cre52, Dal55a, Dau54, Der52, Fit55, Fit56, Fra52, FM56a, Ger53, Ger51, Gul52, Hay50, Hoa51, Jon54, Lin52, Lof52, Lor53, Mur56, Ric53b, Rot58, Sag53, Sar57, Sie57, Spu52, Ste55b, Tep55, Tep52, Vol53, Wal51b, Was58, Woy54, Wri52, ADG^{+}52, Ano53b, Fre53, Har51, Mos50a, WW50].

Status [WR55]. Stuart [Ano54-28]. Studies [Gol53]. Study [CPT56]. Summary [Ber53]. Survey [Har58, Jaf54, Sta50, Yat50]. Syllabus [Mah50, Yat50]. System [Car54, Leo54, Mar54, Wic54].

Teaching [Fit55, FM56a, Rot58, Sar57, Spr59, WD50, Was58]. Technical [Mos50a, CCFM53]. Techniques [Kle59]. Test [Ban57, Cur57, Dos57, Kna57]. Tests [Spr59]. Their [Vol53]. Theory [Gil58]. Thinking [Ros50]. Third [Hae51]. Time [Mah58, Pol59]. Tobacco [Dor54]. Tool [Stu50]. Training [Ald53, Bat50, But59, DK50, Fra52, Fre53, Gul52, Mor50b, Nic54, Olm51, Tep52]. Trials [Gre59]. Turkey [Was58]. Two [Cur57, Stu50]. Two-Group [Cur57]. Two-Way [Stu50]. Type [Ano56g, BQ56, CPT56, Ano56g, BQ56].

U.S. [Ano54i, KDOR51, OCMH50]. U.S.S.R. [Hol58, Ull59]. Undergraduate [Fra52, Har51, Tep52]. Unemployment [Dal55a, Ste55a, Ste55b, Tep55]. Union [Ric52]. United [Lin52, Ano50-27, Fit57, Leo52, Lit50, Lit51, Sim56, Wic54]. Univac [Fer56b, Shi55]. Universities [Cha55a, Fit55, Joh53, Spu52]. University [Ano51i, Cha55b, FP51b, Ano54b, Har51]. Upon [Pol59, Vin52]. Urban [Bag52]. Usability [Ber53, Vol53]. Use [Cre52, Der52, Dor54, Was58]. Uses [Ste55b, Tep55]. USSR [Ano52o].

Value [GL54]. Variation [KS58]. Variations [Arm51]. Various [Gul52]. View [Fut58, Jaf54]. Viewpoints [Ric52]. Vining [Gla52]. Vital [Lor53, Sag53].

Wald [Ano51-31, Hot51]. Way [Stu50]. Weight [Sma54]. Weiss [K.55]. Western [Ano54b, Spu52]. Weyr [Ano51j]. Who [Yat50]. William [Ric59]. Wives [Mor50a]. Work [Ano52z]. Workers [Ano59-29, Gol53]. Works [Fit56]. World [Abr51].

Year [Ano51-29]. Yearbook [Hug55]. You’re [Ano50-30].

References

Abraham:1951:DWI

[Abr51] William I. Abraham. The distribution of world income. The American Statistician, 5(2):6–8, April/May 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

          jstor.org/stable/2681633.

Alpert:1952:RCS

[ADG^{+}52] Harry Alpert, Paul M. Densen, Anne E. Geddes, Margaret Jarman Hagood, Morris H. Hansen, Edward B. Olds, and Conrad Taeuber. Report of the Committee on Social Statistics. The American Statistician, 6(2):1, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681704.

Alderson:1953:STM

[Ald53] Wroe Alderson. Statistical training for marketing research. The American Statistician, 7(1):9–11, February 1953. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682282.

Alpert:1953:RCD

[Alp53] Harry Alpert. The review and coordination of data collecting activities sponsored by the Federal Government. The American Statistician, 7(5):12–14, December 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685716.

Anonymous:1950:AMA

[Ano50a] Anonymous. 110th Annual Meeting of the American Statistical Association. The American Statistician, 4(4):9–18, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685547.

Anonymous:1950:ARS

[Ano50b] Anonymous. Abstracts from the Resolutions of the Second Inter-American Statistical Congress. The American Statistician, 4(2): 18–19, April/May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681408.

Anonymous:1950:AA

[Ano50c] Anonymous. Association activities. The American Statistician, 4 (5):1, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681440.

Anonymous:1950:CNa

[Ano50d] Anonymous. Chapter notes. The American Statistician, 4(1):11–12, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682556.

Anonymous:1950:CNb

[Ano50e] Anonymous. Chapter notes. The American Statistician, 4(2):21–22, April/May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681410.

Anonymous:1950:CNc

[Ano50f] Anonymous. Chapter notes. The American Statistician, 4(3):22, June/July 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682344.

Anonymous:1950:CNd

[Ano50g] Anonymous. Chapter notes. The American Statistician, 4(4):30, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685555.

Anonymous:1950:CNe

[Ano50h] Anonymous. Chapter notes. The American Statistician, 4(5):22, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681452.

Anonymous:1950:CAR

[Ano50i] Anonymous. Committees and Association representatives for 1950. The American Statistician, 4(2):5, April/May 1950. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681401.

Anonymous:1950:FMa

[Ano50j] Anonymous. Front matter. The American Statistician, 4(1):??, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682549.

Anonymous:1950:FMb

[Ano50k] Anonymous. Front matter. The American Statistician, 4(2):4, April/May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681398.

Anonymous:1950:FMc

[Ano50l] Anonymous. Front matter. The American Statistician, 4(3):4, June/July 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682334.

Anonymous:1950:FMd

[Ano50m] Anonymous. Front matter. The American Statistician, 4(4):??, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685542.

Anonymous:1950:FMe

[Ano50n] Anonymous. Front matter. The American Statistician, 4(5):??, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681439.

Anonymous:1950:ISCb

[Ano50o] Anonymous. The Illinois Statistics Conference. The American Statistician, 4(4):5, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685545.

Anonymous:1950:ISCa

[Ano50p] Anonymous. The Institute of Statistics at Chapel Hill. The American Statistician, 4(1):13, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682557.

Anonymous:1950:Na

[Ano50q] Anonymous. News. The American Statistician, 4(1):2–16, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682551.

Anonymous:1950:Nb

[Ano50r] Anonymous. News. The American Statistician, 4(2):2–15, April/ May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681400.

Anonymous:1950:Nc

[Ano50s] Anonymous. News. The American Statistician, 4(3):2–18, June/ July 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682336.

Anonymous:1950:Nd

[Ano50t] Anonymous. News. The American Statistician, 4(4):2–29, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685544.

Anonymous:1950:Ne

[Ano50u] Anonymous. News. The American Statistician, 4(5):2–20, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681441.

Anonymous:1950:NAMa

[Ano50v] Anonymous. News about members. The American Statistician, 4(1):13, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682558.

Anonymous:1950:NAMb

[Ano50w] Anonymous. News about members. The American Statistician, 4(2):22, April/May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681411.

Anonymous:1950:NAMc

[Ano50x] Anonymous. News about members. The American Statistician, 4(3):21, June/July 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682343.

Anonymous:1950:NAMd

[Ano50y] Anonymous. News about members. The American Statistician, 4 (4):29, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685554.

Anonymous:1950:NAMe

[Ano50z] Anonymous. News about members. The American Statistician, 4 (5):10, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681447.

Anonymous:1950:PUN

[Ano50-27] Anonymous. The program of the United Nations Educational Scientific and Cultural Organization. Abstracts from the Draft Program Relating to Social Sciences. The American Statistician, 4(2):16–19, April/May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681406.

Anonymous:1950:RCE

[Ano50-28] Anonymous. Report of the Committee on Elections. The American Statistician, 4(4):1, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685543.

Anonymous:1950:SIA

[Ano50-29] Anonymous. Second Inter-American Statistical Congress. The American Statistician, 4(2):17, April/May 1950. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681407.

Anonymous:1950:YGP

[Ano50-30] Anonymous. So you’re going to present a paper. The American Statistician, 4(4):6–8, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685546.

Anonymous:1950:SQC

[Ano50-31] Anonymous. Statistics and quality control in engineering courses. The American Statistician, 4(3):17, June/July 1950. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682341.

Anonymous:1951:AMA

[Ano51a] Anonymous. 111th Annual Meeting of the American Statistical Association. The American Statistician, 5(4):6–14, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682256.

Anonymous:1951:RCI

[Ano51b] Anonymous. 1950 report of the Committee on Institutional Memberships. The American Statistician, 5(2):1–18, April/May 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681631.

Anonymous:1951:AA

[Ano51c] Anonymous. Association activities. The American Statistician, 5 (3):10–15, June/July 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685986.

Anonymous:1951:CNa

[Ano51d] Anonymous. Chapter notes. The American Statistician, 5(1):22, February 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685923.

Anonymous:1951:CNb

[Ano51e] Anonymous. Chapter notes. The American Statistician, 5(3):20–21, June/July 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685989.

Anonymous:1951:CNc

[Ano51f] Anonymous. Chapter notes. The American Statistician, 5(4):5, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682255.

Anonymous:1951:CNd

[Ano51g] Anonymous. Chapter notes. The American Statistician, 5(5):18, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681868.

Anonymous:1951:CAR

[Ano51h] Anonymous. Committees and Association representatives for 1951. The American Statistician, 5(2):13, April/May 1951. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681636.

Anonymous:1951:CIC

[Ano51i] Anonymous. Courses and instructors in the curriculum of the Faculty of Statistical, Demographic and Actuarial Sciences at the University of Rome. The American Statistician, 5(4):25, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682263.

Anonymous:1951:FW

[Ano51j] Anonymous. Frantisek Weyr, 1879–1951. The American Statistician, 5(5):3, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681859.

Anonymous:1951:FMa

[Ano51k] Anonymous. Front matter. The American Statistician, 5(1):4, February 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685913.

Anonymous:1951:FMb

[Ano51l] Anonymous. Front matter. The American Statistician, 5(2):??, April/May 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681630.

Anonymous:1951:FMc

[Ano51m] Anonymous. Front matter. The American Statistician, 5(3):??, June/July 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685981.

Anonymous:1951:FMd

[Ano51n] Anonymous. Front matter. The American Statistician, 5(4):15, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682252.

Anonymous:1951:FMe

[Ano51o] Anonymous. Front matter. The American Statistician, 5(5):??, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681856.

Anonymous:1951:LE

[Ano51p] Anonymous. Letters to the Editor. The American Statistician, 5(4): 1, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682253.

Anonymous:1951:MN

[Ano51q] Anonymous. More news. The American Statistician, 5(3):14–19, June/July 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685987.

Anonymous:1951:Na

[Ano51r] Anonymous. News. The American Statistician, 5(1):2–13, February 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685915.

Anonymous:1951:Nb

[Ano51s] Anonymous. News. The American Statistician, 5(2):2–5, April/ May 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681632.

Anonymous:1951:Nc

[Ano51t] Anonymous. News. The American Statistician, 5(3):2–5, June/ July 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685983.

Anonymous:1951:Nd

[Ano51u] Anonymous. News. The American Statistician, 5(4):2–27, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682254.

Anonymous:1951:Ne

[Ano51v] Anonymous. News. The American Statistician, 5(5):2–7, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681858.

Anonymous:1951:NAMa

[Ano51w] Anonymous. News about members. The American Statistician, 5(1):13, February 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685918.

Anonymous:1951:NAMb

[Ano51x] Anonymous. News about members. The American Statistician, 5(2):18, April/May 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681640.

Anonymous:1951:NAMc

[Ano51y] Anonymous. News about members. The American Statistician, 5(3):22, June/July 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685990.

Anonymous:1951:NAMd

[Ano51z] Anonymous. News about members. The American Statistician, 5(4):29–30, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682266.

Anonymous:1951:NAMe

[Ano51-27] Anonymous. News about members. The American Statistician, 5(5):17, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681867.

Anonymous:1951:PES

[Ano51-28] Anonymous. Program of the Econometric Society. The American Statistician, 5(5):14–16, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681865.

Anonymous:1951:RCC

[Ano51-29] Anonymous. Report of the Committee on Committees for the year 1950. The American Statistician, 5(4):19–20, October 1951. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682259.

Anonymous:1951:RCE

[Ano51-30] Anonymous. Report of the Committee on Elections. The American Statistician, 5(3):1–4, June/July 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685982.

Anonymous:1951:RHA

[Ano51-31] Anonymous. Resolution in honor of Abraham Wald. The American Statistician, 5(1):19, February 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685921.

Anonymous:1952:AMA

[Ano52a] Anonymous. 112th Annual Meeting of the American Statistical Association. The American Statistician, 6(4):18–25, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682700.

Anonymous:1952:AFI

[Ano52b] Anonymous. Average of forecasts of industrial production index made by leading economists for J. A. Livingston. The American Statistician, 6(2):11, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681710.

Anonymous:1952:CNa

[Ano52c] Anonymous. Chapter notes. The American Statistician, 6(1):22, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682758.

Anonymous:1952:CNb

[Ano52d] Anonymous. Chapter notes. The American Statistician, 6(2):30, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681720.

Anonymous:1952:CNc

[Ano52e] Anonymous. Chapter notes. The American Statistician, 6(3):34, June/July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685980.

Anonymous:1952:CNd

[Ano52f] Anonymous. Chapter notes. The American Statistician, 6(4):30, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682704.

Anonymous:1952:CNe

[Ano52g] Anonymous. Chapter notes. The American Statistician, 6(5):18, December 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681760.

Anonymous:1952:CAR

[Ano52h] Anonymous. Committees and Association representatives for 1952. The American Statistician, 6(2):8–9, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681707.

Anonymous:1952:CAS

[Ano52i] Anonymous. Constitution of the American Statistical Association. The American Statistician, 6(3):23–29, June/July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685976.

Anonymous:1952:FMa

[Ano52j] Anonymous. Front matter. The American Statistician, 6(1):??, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682745.

Anonymous:1952:FMb

[Ano52k] Anonymous. Front matter. The American Statistician, 6(2):5, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681703.

Anonymous:1952:FMc

[Ano52l] Anonymous. Front matter. The American Statistician, 6(3):1, June/July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685969.

Anonymous:1952:FMd

[Ano52m] Anonymous. Front matter. The American Statistician, 6(4):1, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682693.

Anonymous:1952:FMe

[Ano52n] Anonymous. Front matter. The American Statistician, 6(5):1, December 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681750.

Anonymous:1952:MCC

[Ano52o] Anonymous. Methodology Conference of Central Statistical Administration USSR. The American Statistician, 6(3):16–22, June/ July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685975.

Anonymous:1952:MN

[Ano52p] Anonymous. More news. The American Statistician, 6(1):20–22, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682757.

Anonymous:1952:Na

[Ano52q] Anonymous. News. The American Statistician, 6(1):2–4, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682747.

Anonymous:1952:Nb

[Ano52r] Anonymous. News. The American Statistician, 6(2):2–4, April/ May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681705.

Anonymous:1952:Nc

[Ano52s] Anonymous. News. The American Statistician, 6(3):2, June/July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685971.

Anonymous:1952:Nd

[Ano52t] Anonymous. News. The American Statistician, 6(4):2–10, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682694.

Anonymous:1952:Ne

[Ano52u] Anonymous. News. The American Statistician, 6(5):2–3, December 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681751.

Anonymous:1952:NAMa

[Ano52v] Anonymous. News about members. The American Statistician, 6(2):13, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681711.

Anonymous:1952:NAMb

[Ano52w] Anonymous. News about members. The American Statistician, 6(4):29–30, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682703.

Anonymous:1952:PSA

[Ano52x] Anonymous. Philippine Statistical Association. The American Statistician, 6(4):26–27, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682701.

Anonymous:1952:PES

[Ano52y] Anonymous. Program of the Econometric Society. The American Statistician, 6(3):30–31, June/July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685977.

Anonymous:1952:RDA

[Ano52z] Anonymous. Report to the Directors of the American Statistical Association on the Organization and Work of the Advisory Committee on Statistical Policy to the Division of Statistical Standards Bureau of the Budget. The American Statistician, 6(2):10, April/ May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681709.

Anonymous:1952:SC

[Ano52-27] Anonymous. Statistics as a career. The American Statistician, 6(4):6–10, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682696.

Anonymous:1953:AMA

[Ano53a] Anonymous. 113th Annual Meeting of the American Statistical Association. The American Statistician, 7(4):22–27, October 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685766.

Anonymous:1953:AMI

[Ano53b] Anonymous. 16th Annual Meeting of the Institute of Mathematical Statistics. The American Statistician, 7(5):22–24, December 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685719.

Anonymous:1953:CNa

[Ano53c] Anonymous. Chapter notes. The American Statistician, 7(1):17–18, February 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682285.

Anonymous:1953:CNb

[Ano53d] Anonymous. Chapter notes. The American Statistician, 7(2):30, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685639.

Anonymous:1953:CNc

[Ano53e] Anonymous. Chapter notes. The American Statistician, 7(3):25–26, June/July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685498.

Anonymous:1953:CNd

[Ano53f] Anonymous. Chapter notes. The American Statistician, 7(4):33–34, October 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685769.

Anonymous:1953:CNe

[Ano53g] Anonymous. Chapter notes. The American Statistician, 7(5):29–30, December 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685721.

Anonymous:1953:CAR

[Ano53h] Anonymous. Committees and Association representatives for 1953. The American Statistician, 7(2):23–24, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685635.

Anonymous:1953:FMa

[Ano53i] Anonymous. Front matter. The American Statistician, 7(1):??, February 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682276.

Anonymous:1953:FMb

[Ano53j] Anonymous. Front matter. The American Statistician, 7(2):??, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685626.

Anonymous:1953:FMc

[Ano53k] Anonymous. Front matter. The American Statistician, 7(3):24, June/July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685488.

Anonymous:1953:FMd

[Ano53l] Anonymous. Front matter. The American Statistician, 7(4):1, October 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685760.

Anonymous:1953:FMe

[Ano53m] Anonymous. Front matter. The American Statistician, 7(5):1, December 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685712.

Anonymous:1953:Na

[Ano53n] Anonymous. News. The American Statistician, 7(1):2, February 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682278.

Anonymous:1953:Nb

[Ano53o] Anonymous. News. The American Statistician, 7(2):2–4, April/ May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685628.

Anonymous:1953:Nc

[Ano53p] Anonymous. News. The American Statistician, 7(3):2–3, June/ July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685490.

Anonymous:1953:Nd

[Ano53q] Anonymous. News. The American Statistician, 7(4):2, October 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685761.

Anonymous:1953:Ne

[Ano53r] Anonymous. News. The American Statistician, 7(5):2–3, December 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685713.

Anonymous:1953:NAMa

[Ano53s] Anonymous. News about members. The American Statistician, 7(1):16–17, February 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682284.

Anonymous:1953:NAMb

[Ano53t] Anonymous. News about members. The American Statistician, 7(2):28–29, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685638.

Anonymous:1953:NAMc

[Ano53u] Anonymous. News about members. The American Statistician, 7 (3):21–23, June/July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685497.

Anonymous:1953:NAMd

[Ano53v] Anonymous. News about members. The American Statistician, 7(4):29–32, October 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685768.

Anonymous:1953:NAMe

[Ano53w] Anonymous. News about members. The American Statistician, 7(5):25–28, December 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685720.

Anonymous:1953:OLM

[Ano53x] Anonymous. Open letter to the membership. The American Statistician, 7(3):1, June/July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685489.

Anonymous:1953:RCE

[Ano53y] Anonymous. Report of the Committee on Elections. The American Statistician, 7(2):1, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685627.

Anonymous:1954:AMA

[Ano54a] Anonymous. 114th Annual Meeting of the American Statistical Association. The American Statistician, 8(3):24–32, June/July 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685459.

Anonymous:1954:ASA

[Ano54b] Anonymous. American Statistical Association Western Regional meeting: University of California, Berkeley, December 27–31, 1954. The American Statistician, 8(4):24–27, October 1954. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682056.

Anonymous:1954:CNa

[Ano54c] Anonymous. Chapter notes. The American Statistician, 8(1):21–22, February 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681664.

Anonymous:1954:CNb

[Ano54d] Anonymous. Chapter notes. The American Statistician, 8(2):25–26, April/May 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686021.

Anonymous:1954:CNc

[Ano54e] Anonymous. Chapter notes. The American Statistician, 8(3):37–38, June/July 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685463.

Anonymous:1954:CNd

[Ano54f] Anonymous. Chapter notes. The American Statistician, 8(4):40–42, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682062.

Anonymous:1954:CNe

[Ano54g] Anonymous. Chapter notes. The American Statistician, 8(5):25–26, December 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681550.

Anonymous:1954:CAR

[Ano54h] Anonymous. Committees and Association representatives for 1954. The American Statistician, 8(2):20–21, April/May 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2686018.

Anonymous:1954:CPA

[Ano54i] Anonymous. Cooperative placement arranged with U.S. employment service. The American Statistician, 8(4):1, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682049.

Anonymous:1954:ENF

[Ano54j] Anonymous. Election of new Fellows. The American Statistician, 8(1):17–18, February 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681662.

Anonymous:1954:FMa

[Ano54k] Anonymous. Front matter. The American Statistician, 8(1):??, February 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681654.

Anonymous:1954:FMb

[Ano54l] Anonymous. Front matter. The American Statistician, 8(2):24, April/May 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686012.

Anonymous:1954:FMc

[Ano54m] Anonymous. Front matter. The American Statistician, 8(3):??, June/July 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685451.

Anonymous:1954:FMd

[Ano54n] Anonymous. Front matter. The American Statistician, 8(4):35, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682048.

Anonymous:1954:FMe

[Ano54o] Anonymous. Front matter. The American Statistician, 8(5):1, December 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681542.

Anonymous:1954:NSS

[Ano54p] Anonymous. New section on social statistics. The American Statistician, 8(1):15–16, February 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681661.

Anonymous:1954:Na

[Ano54q] Anonymous. News. The American Statistician, 8(1):2–3, February 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681656.

Anonymous:1954:Nb

[Ano54r] Anonymous. News. The American Statistician, 8(2):2–4, April/ May 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686014.

Anonymous:1954:Nc

[Ano54s] Anonymous. News. The American Statistician, 8(3):2–3, June/ July 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685453.

Anonymous:1954:Nd

[Ano54t] Anonymous. News. The American Statistician, 8(4):2–4, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682050.

Anonymous:1954:Ne

[Ano54u] Anonymous. News. The American Statistician, 8(5):2–22, December 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681543.

Anonymous:1954:NAMa

[Ano54v] Anonymous. News about members. The American Statistician, 8(1):19–20, February 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681663.

Anonymous:1954:NAMb

[Ano54w] Anonymous. News about members. The American Statistician, 8(2):22–23, April/May 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686020.

Anonymous:1954:NAMc

[Ano54x] Anonymous. News about members. The American Statistician, 8(3):36, June/July 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685462.

Anonymous:1954:NAMd

[Ano54y] Anonymous. News about members. The American Statistician, 8(4):36–37, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682059.

Anonymous:1954:NAMe

[Ano54z] Anonymous. News about members. The American Statistician, 8(5):23–24, December 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681549.

Anonymous:1954:POA

[Ano54-27] Anonymous. Permanent organization for the Advisory Committee on Statistical Policy to the Office of Statistical Standards, Bureau of the Budget. The American Statistician, 8(3):33, June/July 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685460.

Anonymous:1954:RHS

[Ano54-28] Anonymous. Resolution honoring Stuart A. Rice. The American Statistician, 8(4):38–39, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682060.

Anonymous:1955:BMa

[Ano55a] Anonymous. Back matter. The American Statistician, 9(1):??, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682356.

Anonymous:1955:BMb

[Ano55b] Anonymous. Back matter. The American Statistician, 9(3):28–29, June 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681463.

Anonymous:1955:BMc

[Ano55c] Anonymous. Back matter. The American Statistician, 9(4):??, October 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682121.

Anonymous:1955:CNa

[Ano55d] Anonymous. Chapter notes. The American Statistician, 9(1):28–29, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682355.

Anonymous:1955:CNb

[Ano55e] Anonymous. Chapter notes. The American Statistician, 9(2):29–30, April 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681974.

Anonymous:1955:CNc

[Ano55f] Anonymous. Chapter notes. The American Statistician, 9(3):24–27, June 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681462.

Anonymous:1955:CNd

[Ano55g] Anonymous. Chapter notes. The American Statistician, 9(4):34–36, October 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682120.

Anonymous:1955:CNe

[Ano55h] Anonymous. Chapter notes. The American Statistician, 9(5):25–26, December 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685506.

Anonymous:1955:FMa

[Ano55i] Anonymous. Front matter. The American Statistician, 9(1):1–18, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682345.

Anonymous:1955:FMb

[Ano55j] Anonymous. Front matter. The American Statistician, 9(2):1–26, April 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681966.

Anonymous:1955:FMc

[Ano55k] Anonymous. Front matter. The American Statistician, 9(3):??, June 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681453.

Anonymous:1955:FMd

[Ano55l] Anonymous. Front matter. The American Statistician, 9(4):1–27, October 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682112.

Anonymous:1955:FMe

[Ano55m] Anonymous. Front matter. The American Statistician, 9(5):1–22, December 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685499.

Anonymous:1955:Na

[Ano55n] Anonymous. News. The American Statistician, 9(1):2–4, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682346.

Anonymous:1955:Nb

[Ano55o] Anonymous. News. The American Statistician, 9(2):2–6, April 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681967.

Anonymous:1955:Nc

[Ano55p] Anonymous. News. The American Statistician, 9(3):2–3, June 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681455.

Anonymous:1955:Nd

[Ano55q] Anonymous. News. The American Statistician, 9(4):2–5, October 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682113.

Anonymous:1955:Ne

[Ano55r] Anonymous. News. The American Statistician, 9(5):2–11, December 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685500.

Anonymous:1955:NAMa

[Ano55s] Anonymous. News about members. The American Statistician, 9(1):26–27, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682354.

Anonymous:1955:NAMb

[Ano55t] Anonymous. News about members. The American Statistician, 9 (2):27–28, April 1955. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681973.

Anonymous:1955:NAMc

[Ano55u] Anonymous. News about members. The American Statistician, 9 (3):23–24, June 1955. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681461.

Anonymous:1955:NAMd

[Ano55v] Anonymous. News about members. The American Statistician, 9(4):31–33, October 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682119.

Anonymous:1955:NAMe

[Ano55w] Anonymous. News about members. The American Statistician, 9(5):23–24, December 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685505.

Anonymous:1955:PAM

[Ano55x] Anonymous. Program of the 115th Annual Meeting of the American Statistical Association. The American Statistician, 9(4):18–26, October 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682117.

Anonymous:1956:BM

[Ano56a] Anonymous. Back matter. The American Statistician, 10(3):30, June 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681780.

Anonymous:1956:CNa

[Ano56b] Anonymous. Chapter notes. The American Statistician, 10(1):20– 22, February 1956. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681472.

Anonymous:1956:CNb

[Ano56c] Anonymous. Chapter notes. The American Statistician, 10(2):20– 22, April 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681603.

Anonymous:1956:CNc

[Ano56d] Anonymous. Chapter notes. The American Statistician, 10(3):27– 29, June 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681779.

Anonymous:1956:CNd

[Ano56e] Anonymous. Chapter notes. The American Statistician, 10(4):27– 29, October 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681329.

Anonymous:1956:CNe

[Ano56f] Anonymous. Chapter notes. The American Statistician, 10(5):25– 26, December 1956. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681702.

Anonymous:1956:CMR

[Ano56g] Anonymous. Correction: Multiple regression and correlation analysis on the IBM Type 701 and Type 704 electronic data processing machines. The American Statistician, 10(3):24, June 1956. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681777. See [BQ56].

Anonymous:1956:FMa

[Ano56h] Anonymous. Front matter. The American Statistician, 10(1):1–18, February 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681464.

Anonymous:1956:FMb

[Ano56i] Anonymous. Front matter. The American Statistician, 10(2):1–19, April 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681594.

Anonymous:1956:FMc

[Ano56j] Anonymous. Front matter. The American Statistician, 10(3):1–22, June 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681770.

Anonymous:1956:FMd

[Ano56k] Anonymous. Front matter. The American Statistician, 10(4):19, October 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681320.

Anonymous:1956:FMe

[Ano56l] Anonymous. Front matter. The American Statistician, 10(5):1–22, December 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681696.

Anonymous:1956:Na

[Ano56m] Anonymous. News. The American Statistician, 10(1):2–3, February 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681465.

Anonymous:1956:Nb

[Ano56n] Anonymous. News. The American Statistician, 10(2):2–18, April 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681595.

Anonymous:1956:Nc

[Ano56o] Anonymous. News. The American Statistician, 10(3):2–25, June 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681771.

Anonymous:1956:Nd

[Ano56p] Anonymous. News. The American Statistician, 10(4):2–4, October 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681322.

Anonymous:1956:Ne

[Ano56q] Anonymous. News. The American Statistician, 10(5):2–7, December 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681697.

Anonymous:1956:NAMa

[Ano56r] Anonymous. News about members. The American Statistician, 10(1):19–20, February 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681471.

Anonymous:1956:NAMb

[Ano56s] Anonymous. News about members. The American Statistician, 10(2):17–18, April 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681602.

Anonymous:1956:NAMc

[Ano56t] Anonymous. News about members. The American Statistician, 10(3):26–27, June 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681778.

Anonymous:1956:NAMd

[Ano56u] Anonymous. News about members. The American Statistician, 10(4):24–30, October 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681328.

Anonymous:1956:NAMe

[Ano56v] Anonymous. News about members. The American Statistician, 10(5):23–24, December 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681701.

Anonymous:1956:PAM

[Ano56w] Anonymous. Program of the 116th Annual Meeting of the American Statistical Association. The American Statistician, 10(3):14–21, June 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681775.

Anonymous:1956:RAH

[Ano56x] Anonymous. Report of the Ad Hoc Committee to Explore Opinion on Standards. The American Statistician, 10(1):13–16, February 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681469.

Anonymous:1957:CNa

[Ano57a] Anonymous. Chapter notes. The American Statistician, 11(1):24– 26, February 1957. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681855.

Anonymous:1957:CNb

[Ano57b] Anonymous. Chapter notes. The American Statistician, 11(2):27– 29, April 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681769.

Anonymous:1957:CNc

[Ano57c] Anonymous. Chapter notes. The American Statistician, 11(3):36– 38, June 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685683.

Anonymous:1957:CNd

[Ano57d] Anonymous. Chapter notes. The American Statistician, 11(4):27– 30, October 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681593.

Anonymous:1957:CNe

[Ano57e] Anonymous. Chapter notes. The American Statistician, 11(5):29– 30, December 1957. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682569.

Anonymous:1957:FMa

[Ano57f] Anonymous. Front matter. The American Statistician, 11(1):1–23, February 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681848.

Anonymous:1957:FMb

[Ano57g] Anonymous. Front matter. The American Statistician, 11(2):1–26, April 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681761.

Anonymous:1957:FMc

[Ano57h] Anonymous. Front matter. The American Statistician, 11(3):1–34, June 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685676.

Anonymous:1957:FMd

[Ano57i] Anonymous. Front matter. The American Statistician, 11(4):1–23, October 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681582.

Anonymous:1957:FMe

[Ano57j] Anonymous. Front matter. The American Statistician, 11(5):1–22, December 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682562.

Anonymous:1957:Na

[Ano57k] Anonymous. News. The American Statistician, 11(1):2–10, February 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681849.

Anonymous:1957:Nb

[Ano57l] Anonymous. News. The American Statistician, 11(2):2–4, April 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681762.

Anonymous:1957:Nc

[Ano57m] Anonymous. News. The American Statistician, 11(3):2–5, June 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685677.

Anonymous:1957:Nd

[Ano57n] Anonymous. News. The American Statistician, 11(4):2–5, October 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681583.

Anonymous:1957:Ne

[Ano57o] Anonymous. News. The American Statistician, 11(5):2–6, December 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682563.

Anonymous:1957:NAMa

[Ano57p] Anonymous. News about members. The American Statistician, 11(1):21–22, February 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681854.

Anonymous:1957:NAMb

[Ano57q] Anonymous. News about members. The American Statistician, 11(2):24–25, April 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681768.

Anonymous:1957:NAMc

[Ano57r] Anonymous. News about members. The American Statistician, 11(3):35–36, June 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685682.

Anonymous:1957:NAMd

[Ano57s] Anonymous. News about members. The American Statistician, 11(4):25–27, October 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681592.

Anonymous:1957:NAMe

[Ano57t] Anonymous. News about members. The American Statistician, 11(5):27–28, December 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682568.

Anonymous:1957:PAM

[Ano57u] Anonymous. Program of the 117th Annual Meeting of the American Statistical Association. The American Statistician, 11(3):20–30, June 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685680.

Anonymous:1958:BMa

[Ano58a] Anonymous. Back matter. The American Statistician, 12(2):30, April 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681516.

Anonymous:1958:BMb

[Ano58b] Anonymous. Back matter. The American Statistician, 12(3):30, June 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682548.

Anonymous:1958:CNa

[Ano58c] Anonymous. Chapter notes. The American Statistician, 12(1):32– 34, February 1958. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682676.

Anonymous:1958:CNb

[Ano58d] Anonymous. Chapter notes. The American Statistician, 12(2):27– 28, April 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681515.

Anonymous:1958:CNc

[Ano58e] Anonymous. Chapter notes. The American Statistician, 12(3):27– 29, June 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682547.

Anonymous:1958:CNd

[Ano58f] Anonymous. Chapter notes. The American Statistician, 12(4):39– 40, October 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681790.

Anonymous:1958:CNe

[Ano58g] Anonymous. Chapter notes. The American Statistician, 12(5):20– 21, December 1958. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682275.

Anonymous:1958:FMa

[Ano58h] Anonymous. Front matter. The American Statistician, 12(1):1–27, February 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682666.

Anonymous:1958:FMb

[Ano58i] Anonymous. Front matter. The American Statistician, 12(2):19–24, April 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681506.

Anonymous:1958:FMc

[Ano58j] Anonymous. Front matter. The American Statistician, 12(3):1–20, June 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682540.

Anonymous:1958:FMd

[Ano58k] Anonymous. Front matter. The American Statistician, 12(4):23, October 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681781.

Anonymous:1958:FMe

[Ano58l] Anonymous. Front matter. The American Statistician, 12(5):15, December 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682267.

Anonymous:1958:ISA

[Ano58m] Anonymous. International statistical activities. The American Statistician, 12(5):6–18, December 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682270.

Anonymous:1958:Na

[Ano58n] Anonymous. News. The American Statistician, 12(1):2–7, February 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682667.

Anonymous:1958:Nb

[Ano58o] Anonymous. News. The American Statistician, 12(2):2–7, April 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681508.

Anonymous:1958:Nc

[Ano58p] Anonymous. News. The American Statistician, 12(3):2–17, June 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682541.

Anonymous:1958:Nd

[Ano58q] Anonymous. News. The American Statistician, 12(4):2–8, October 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681783.

Anonymous:1958:Ne

[Ano58r] Anonymous. News. The American Statistician, 12(5):2–5, December 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682269.

Anonymous:1958:NAMa

[Ano58s] Anonymous. News about members. The American Statistician, 12(1):31–32, February 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682675.

Anonymous:1958:NAMb

[Ano58t] Anonymous. News about members. The American Statistician, 12(2):25–26, April 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681514.

Anonymous:1958:NAMc

[Ano58u] Anonymous. News about members. The American Statistician, 12(3):25–26, June 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682546.

Anonymous:1958:NAMd

[Ano58v] Anonymous. News about members. The American Statistician, 12(4):36–39, October 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681789.

Anonymous:1958:NAMe

[Ano58w] Anonymous. News about members. The American Statistician, 12(5):19–20, December 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682274.

Anonymous:1958:PAM

[Ano58x] Anonymous. Program of the 118th Annual Meeting of the American Statistical Association. The American Statistician, 12(4):26–34, October 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681788.

Anonymous:1959:BMa

[Ano59a] Anonymous. Back matter. The American Statistician, 13(2):29–30, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682317.

Anonymous:1959:BMb

[Ano59b] Anonymous. Back matter. The American Statistician, 13(5):41, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682437.

Anonymous:1959:CNa

[Ano59c] Anonymous. Chapter notes. The American Statistician, 13(1):27– 29, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685866.

Anonymous:1959:CNb

[Ano59d] Anonymous. Chapter notes. The American Statistician, 13(2):27– 28, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682316.

Anonymous:1959:CNc

[Ano59e] Anonymous. Chapter notes. The American Statistician, 13(3):31– 33, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681372.

Anonymous:1959:CNd

[Ano59f] Anonymous. Chapter notes. The American Statistician, 13(4):41– 44, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685758.

Anonymous:1959:CNe

[Ano59g] Anonymous. Chapter notes. The American Statistician, 13(5):39– 40, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682436.

Anonymous:1959:FMa

[Ano59h] Anonymous. Front matter. The American Statistician, 13(1):8–25, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685855.

Anonymous:1959:FMb

[Ano59i] Anonymous. Front matter. The American Statistician, 13(2):22–25, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682306.

Anonymous:1959:FMc

[Ano59j] Anonymous. Front matter. The American Statistician, 13(3):26, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681359.

Anonymous:1959:FMd

[Ano59k] Anonymous. Front matter. The American Statistician, 13(4):16–30, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685746.

Anonymous:1959:FMe

[Ano59l] Anonymous. Front matter. The American Statistician, 13(5):18–28, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682424.

Anonymous:1959:ISAa

[Ano59m] Anonymous. International statistical activities. The American Statistician, 13(1):9–10, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685858.

Anonymous:1959:ISAb

[Ano59n] Anonymous. International statistical activities. The American Statistician, 13(2):7–8, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682309.

Anonymous:1959:ISAc

[Ano59o] Anonymous. International statistical activities. The American Statistician, 13(3):5, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681362.

Anonymous:1959:ISAd

[Ano59p] Anonymous. International statistical activities. The American Statistician, 13(4):7–8, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685749.

Anonymous:1959:ISAe

[Ano59q] Anonymous. International statistical activities. The American Statistician, 13(5):7–8, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682427.

Anonymous:1959:Na

[Ano59r] Anonymous. News. The American Statistician, 13(1):2–7, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685857.

Anonymous:1959:Nb

[Ano59s] Anonymous. News. The American Statistician, 13(2):2–6, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682308.

Anonymous:1959:Nc

[Ano59t] Anonymous. News. The American Statistician, 13(3):2–4, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681361.

Anonymous:1959:Nd

[Ano59u] Anonymous. News. The American Statistician, 13(4):2–6, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685748.

Anonymous:1959:Ne

[Ano59v] Anonymous. News. The American Statistician, 13(5):2–6, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682426.

Anonymous:1959:NAMa

[Ano59w] Anonymous. News about members. The American Statistician, 13(1):26–27, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685865.

Anonymous:1959:NAMb

[Ano59x] Anonymous. News about members. The American Statistician, 13 (2):26, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682315.

Anonymous:1959:NAMc

[Ano59y] Anonymous. News about members. The American Statistician, 13(3):30–31, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681371.

Anonymous:1959:NAMd

[Ano59z] Anonymous. News about members. The American Statistician, 13(4):39–41, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685757.

Anonymous:1959:NAMe

[Ano59-27] Anonymous. News about members. The American Statistician, 13(5):38–39, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682435.

Anonymous:1959:PAM

[Ano59-28] Anonymous. Program of the 119th Annual Meeting of the American Statistical Association. The American Statistician, 13(4):21–29, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685754.

Anonymous:1959:RAF

[Ano59-29] Anonymous. Recommendations on availability of federal statistical materials to nongovernmental research workers. The American Statistician, 13(4):15–37, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685752.

Armstrong:1950:ECSa

[Arm50a] Charles E. Armstrong. The effects of calendar shifts on series of monthly data. The American Statistician, 4(4):20–21, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685549.

Armstrong:1950:ECSb

[Arm50b] Charles E. Armstrong. The effects of calendar shifts on series of monthly data. The American Statistician, 4(5):11–13, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681448.

Armstrong:1951:SVC

[Arm51] Charles Armstrong. Seasonal variations and calendar influences. The American Statistician, 5(5):10, December 1951. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681862.

Bagley:1952:UML

[Bag52] Eleanor S. Bagley. Urban mortgage loan statistics in One Life Company. The American Statistician, 6(4):11–14, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682697.

Bangdiwala:1957:GFT

[Ban57] I. S. Bangdiwala. A general formula for the t test. The American Statistician, 11(4):21, October 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681589. See comments [Dwy58].

Bancroft:1958:SC

[Ban58] T. A. Bancroft. From the standpoint of colleges. The American Statistician, 12(1):20–21, February 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682670.

Baten:1950:AOC

[Bat50] William Dowell Baten. Another opinion concerning statistical training below the college level. The American Statistician, 4(4):24, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685551.

Bowman:1959:FSAb

[BD59] Raymond T. Bowman and Samuel J. Dennis. Federal statistical activities. The American Statistician, 13(4):9–13, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685750.

Bowman:1957:FSA

[BDG57] Raymond T. Bowman, Wayne F. Daugherty, and Edwin D. Goldfield. Federal statistical activities. The American Statistician, 11 (1):12–14, February 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681850.

Berkeley:1950:EHP

[Ber50] E. C. Berkeley. Eight hundred people interested in mechanical brains. The American Statistician, 4(3):11–12, June/July 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682338.

Bergson:1953:RUS

[Ber53] Abram Bergson. Reliability and usability of Soviet statistics: A summary appraisal. The American Statistician, 7(3):13–16, June/ July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685494.

Bowman:1959:FSAa

[BFR59] Raymond T. Bowman, David L. Futransky, and Israel Rubin. Federal statistical activities. The American Statistician, 13(1):11–23, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685859.

Bowman:1955:FSA

[BGS^{+}55] Raymond T. Bowman, Edwin D. Goldfield, Julius Shiskin, Howard G. Brunsman, T. D. Woolsey, Laughling F. McHugh, John W. Lehman, and William Lerner. Federal statistical activities. The American Statistician, 9(5):4–7, December 1955. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685501.

Brown:1951:SD

[BL51] Theodore H. Brown and Dickson H. Leavens. A sampling device. The American Statistician, 5(4):16–17, October 1951. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682257.

Bolling:1956:RSS

[Bol56] Richard Bolling. The role of statistics in shaping economic policy. The American Statistician, 10(3):7–9, June 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681773.

Bonnell:1950:PEL

[Bon50] L. S. Bonnell. Preparation of effective lantern slides. The American Statistician, 4(4):25–26, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685552.

Bowman:1952:RPS

[Bow52] Raymond T. Bowman. Report on Production Scheduling Section. The American Statistician, 6(5):12–13, December 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681756.

Bowman:1957:IFS

[Bow57] Raymond T. Bowman. Improvement of federal statistics. The American Statistician, 11(2):18–21, April 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681766.

Beckman:1956:MRC

[BQ56] F. S. Beckman and D. A. Quarles, Jr. Multiple regression and correlation analysis on the IBM Type 701 and Type 704 electronic data processing machines. The American Statistician, 10(1):6–16, February 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681467.

See correction [Ano56g].

Brown:1952:SHC

[Bro52] Theodore H. Brown. “the statistician and his conscience”. The American Statistician, 6(1):14–18, February 1952. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682755.

Broida:1955:DI

[Bro55] Arthur L. Broida. Diffusion indexes. The American Statistician, 9(3):7–16, June 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681458.

Bowman:1958:FSA

[BTG^{+}58] Raymond T. Bowman, Conrad Taeuber, Louis Greenberg, Sidney A. Jaffe, Wayne F. Daugherty, and William Lerner. Federal statistical activities. The American Statistician, 12(1):8–12, February 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682668.

Burr:1950:SEP

[Bur50] Irving W. Burr. Statistics for engineers at purdue. The American Statistician, 4(4):19, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685548.

Bursk:1959:CSP

[Bur59] J. Parker Bursk. A classification of sampling problems. The American Statistician, 13(2):19–21, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682313.

Butterbaugh:1959:ETF

[But59] Grant I. Butterbaugh. Educational and training films in statistics. The American Statistician, 13(2):13–17, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682311.

Broida:1952:FSA

[BWW^{+}52] Arthur L. Broida, Samuel Weiss, O. V. Wells, Edward Hollander, Howard Ennes, and Morris B. Ullman. Federal statistical activities. The American Statistician, 6(2):6–7, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681706.

Carter:1954:PCO

[Car54] C. F. Carter. [problems of co-ordination in the Canadian Statistical System]: Comments. The American Statistician, 8(4):22–23, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682055.

Carson:1953:RAAb

[CCD^{+}53] William J. Carson, William G. Cochran, E. Dana Durand, Simon Kuznets, Isador Lubin, William F. Ogburn, Lowell J. Reed, and Samuel S. Wilks. Report of ASA Advisory Committee on Statistical Policy to the Bureau of the Budget. The American Statistician, 7(5):20–21, December 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685718.

Calvin:1959:RCP

[CCD^{+}59] Lyle D. Calvin, William S. Connor, Harold F. Dorn, Bernard G. Greenberg, Morris Hamburg, Frank A. Hanna, Earl E. Houseman, J. Stuart Hunter, John W. Lehman, John Neter, Donald C. Riley, W. Allen Wallis, and Almarin Phillips. Report of the Committee on Publications Policy on the Results of the Publications Questionnaire. The American Statistician, 13(5):29–37, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682434.

Clausen:1954:RAH

[CCF^{+}54] John A. Clausen, Herbert S. Conrad, William W. K. Freeman, Nicholas Hobbs, John W. Hopkins, Alfred N. Watson, and Rensis Likert. Report of the Ad Hoc Committee on Statistical Standards. The American Statistician, 8(3):19–23, June/July 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685458.

Cowden:1953:RTC

[CCFM53] Dudley Cowden, Reavis Cox, Solomon Fabricant, and Bruce D. Mudgett. Report of the Technical Committee of the American Statistical Association. The American Statistician, 7(3):17–20, June/ July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685495.

Copeland:1951:RCF

[CDK^{+}51] Morris Copeland, W. Edwards Deming, Lester S. Kellogg, Frank W. Notestein, and George W. Snedecor. Report of the Committee on Fellows. The American Statistician, 5(1):1, February 1951. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685914.

Carson:1953:RAAa

[CDK^{+}53] William J. Carson, E. Dana Durand, Simon Kuznets, Isador Lubin, William F. Ogburn, Lowell J. Reed, and Samuel S. Wilks. Report of ASA Advisory Committee on Statistical Policy to the Bureau of the Budget. The American Statistician, 7(4):28–33, October 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685767.

Conklin:1958:SEE

[CG58] Maxwell R. Conklin and Owen C. Gretton. Some experience with electronic computers in processing the 1954 Census of Manufactures. The American Statistician, 12(4):19–42, October 1958. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681786.

Chapman:1955:ISC

[Cha55a] H. H. Chapman. Instruction in statistics in the colleges and universities of the South. The American Statistician, 9(2):18–21, April 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681971.

Chapman:1955:PPU

[Cha55b] H. H. Chapman. Problems of providing a university program in statistics. The American Statistician, 9(3):17–22, June 1955. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681459.

Cochran:1952:RCE

[CLE52] William G. Cochran, Irving Lorge, and R. J. Eggert. Report of the Committee on Elections. The American Statistician, 6(3):1, June/ July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685970.

Cole:1959:SC

[Col59] LaMont C. Cole. On simplified computations. The American Statistician, 13(1):20, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685861.

Compton:1950:MPI

[Com50] W. Dean Compton. Map presentation III. Avoiding the “small state” problem. The American Statistician, 4(1):17, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682560.

Court:1952:SS

[Cou52] Andrew T. Court. Statistical standards. The American Statistician, 6(1):6, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682749.

Cox:1952:RDP

[Cox52] Reavis Cox. Report on Distribution Panel. The American Statistician, 6(5):14–15, December 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681758.

Caffrey:1956:ESE

[CPT56] John Caffrey, N. C. Perry, and D. Teichroew. An empirical study of an estimate of skewness for the Pearson Type III distribution. The American Statistician, 10(2):13–14, April 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681599.

Creamer:1952:SCU

[Cre52] Daniel Creamer. Some comments on “The Use and Development of National Income Statistics”. The American Statistician, 6(2):18–19, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681715. See [Der52].

Connor:1953:QAc

[CSC^{+}53] W. S. Connor, I. R. Savage, J. C., E. P. King, and Seymour Banks. Questions and answers. The American Statistician, 7(3):19–20, June/July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685496.

Connor:1952:QA

[CSDP52] W. S. Connor, I. R. Savage, C. Derman, and J. Q. P. Questions and answers. The American Statistician, 6(4):28, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682702.

Connor:1953:QAb

[CSP^{+}53a] W. S. Connor, I. R. Savage, J. Q. P., Morton S. Raff, and Paul R. Rider. Questions and answers. The American Statistician, 7 (2):26–27, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685637.

Connor:1953:QAa

[CSP^{+}53b] W. S. Connor, I. R. Savage, F. Proschan, S. Kullback, and William Kruskal. Questions and answers. The American Statistician, 7 (1):14–15, February 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682283.

Cureton:1957:FNT

[Cur57] Edward E. Cureton. Further note on the two-group t test. The American Statistician, 11(4):21, October 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681591. See comments [Dwy58].

Dale:1955:RME

[Dal55a] D. K. Dale. The reliability and meaning of employment and unemployment statistics: Discussion. The American Statistician, 9 (1):14–15, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682350.

Daly:1955:MAG

[Dal55b] Joseph F. Daly. Meyer Abraham Girshick, 1908–1955. The American Statistician, 9(3):6, June 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681457.

Daly:1957:OPR

[Dal57] Joseph F. Daly. Organizational problems related to large scale statistical computations at Bureau of Census. The American Statistician, 11(2):10–23, April 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681764.

Dauer:1954:LCC

[Dau54] Ernst A. Dauer. Limitations of consumer credit statistics. The American Statistician, 8(3):15–18, June/July 1954. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685457.

Day:1950:SMA

[Day50] Besse B. Day. Statistical methods applied to proximity fuzes for projectiles. The American Statistician, 4(2):6–11, April/May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681402.

Day:1953:SKS

[Day53] Price Day. Sliderule, Kelly, sliderule! The American Statistician, 7(1):7–8, February 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682281.

Day:1954:DEN

[Day54] B. B. Day. The design of experiments in naval engineering. The American Statistician, 8(2):7–23, April/May 1954. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2686016.

Dixon:1954:LE

[DC54] W. J. Dixon and Lee J. Cronbach. Letters to the Editor. The American Statistician, 8(4):39, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682061.

Deming:1958:SPS

[Dem58] W. Edwards Deming. Standards of probability sampling for legal evidence. The American Statistician, 12(1):25–26, February 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682673.

Derksen:1952:UDN

[Der52] J. B. D. Derksen. The use and development of national income statistics. The American Statistician, 6(2):16–18, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681714. See comments

[Cre52].

Dutka:1950:STB

[DK50] S. Dutka and F. Kafka. Statistical training below the college level. The American Statistician, 4(1):6–7, February 1950. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682553.

Drury:1956:FSA

[DKC^{+}56] Robert F. Drury, David L. Kaplan, Herbert S. Conrad, D. A. Williams, Jean L. Pennock, Leon R. Paley, Dudley E. Young, and Donald E. Church. Federal statistical activities. The American Statistician, 10(5):8–13, December 1956. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681698.

Donnahoe:1951:OAN

[Don51] Alan S. Donnahoe. One answer to the “no change” problem. The American Statistician, 5(2):14, April/May 1951. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681637.

Dorman:1950:PNP

[Dor50] Robert O. Dorman. A proposal for a National Placement Service. The American Statistician, 4(1):8, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682554.

Dorn:1951:RAM

[Dor51] Harold F. Dorn. Report of the 1951 Annual Meeting Program Committee. The American Statistician, 5(5):11, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681863.

Dorn:1954:RCL

[Dor54] Harold F. Dorn. The relationship of cancer of the lung and the use of tobacco. The American Statistician, 8(5):7–13, December 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681545.

Dossett:1957:FNT

[Dos57] William F. Dossett. A further note on the t test. The American Statistician, 11(4):21, October 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681590. See comments [Dwy58].

Dungee:1955:QA

[Dun55] Shirley Dungee. Questions and answers. The American Statistician, 9(4):28–30, October 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682118.

Dwyer:1957:BIS

[Dwy57] Paul S. Dwyer. Basic instructions in statistical computations. The American Statistician, 11(2):13–17, April 1957. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681765.

Dwyer:1958:CF

[Dwy58] Paul S. Dwyer. Computational formulas for t^{2}. The American Statistician, 12(3):18–19, June 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www. jstor.org/stable/2682544. See [Ban57, Cur57, Dos57, Kna57].

Elkins:1952:CC

[EBL52] Thomas A. Elkins, Robert W. Burgess, and Ben Lipstein. Changes in the constitution. The American Statistician, 6(4):14, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682698.

Eckler:1953:ECE

[Eck53] A. Ross Eckler. Extent and character of errors in the 1950 Census. The American Statistician, 7(5):15–21, December 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685717.

Edwards:1959:CSC

[Edw59] Bruce Edwards. Constructing simple correlation problems with predetermined answers. The American Statistician, 13(5):25–27, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682433.

Ferber:1956:IAS

[Fer56a] Robert Ferber. On the interpretation of the aggregate savings ratio. The American Statistician, 10(3):10–13, June 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681774.

Ferger:1956:ACU

[Fer56b] Wirth F. Ferger. Admissibility in court of Univac seasonal computations. The American Statistician, 10(1):9–10, February 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681468.

Fitzpatrick:1955:ETS

[Fit55] Paul J. Fitzpatrick. The early teaching of statistics in American colleges and universities. The American Statistician, 9(5):12–18, December 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685503.

Fitzpatrick:1956:SWE

[Fit56] Paul J. Fitzpatrick. Statistical works in early American statistics courses. The American Statistician, 10(5):14–19, December 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681699.

FitzPatrick:1957:SSU

[Fit57] Paul J. FitzPatrick. Statistical societies in the United States in the Nineteenth Century. The American Statistician, 11(5):13–21, December 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682565.

Freund:1955:DF

[FM55] John E. Freund and Irwin Miller. In defense of frequencies. The American Statistician, 9(2):15–27, April 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681970.

Freeman:1956:TIS

[FM56a] Linton C. Freeman and Douglas M. More. Teaching introductory statistics in the liberal arts curriculum. The American Statistician, 10(4):20–23, October 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681327.

Freund:1956:SSA

[FM56b] John E. Freund and Irwin Miller. Some statistical aspects of a problem in scientific philosophy. The American Statistician, 10(2): 7–11, April 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681597.

Farrell:1951:HFH

[FP51a] Helen Farrell and Robert H. Perry. How forecasting helps the Chicago Chapter. The American Statistician, 5(5):16, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681866.

Federici:1951:FSD

[FP51b] Nora Federici and Fernando Pedroni. The Faculty of Statistical, Demographic and Actuarial Sciences at the University of Rome.

The American Statistician, 5(4):22–24, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682262.

Frankel:1952:DUT

[Fra52] Lester R. Frankel. [A desirable undergraduate training program in statistics]: Discussion. The American Statistician, 6(5):11–15, December 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681755.

Freeman:1952:SHC

[Fre52] William W. K. Freeman. [“the statistician and his conscience”]: Discussion. The American Statistician, 6(1):18–20, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682756.

Freund:1953:RCT

[Fre53] J. E. Freund. Report of the Committee on Training in Statistics. The American Statistician, 7(2):24–25, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685636.

Futransky:1958:PVF

[Fut58] David L. Futransky. From the point of view of the Federal Government. The American Statistician, 12(1):21–23, February 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682671.

Galt:1952:FSA

[Gal52] Lowell T. Galt. Federal statistical activities. The American Statistician, 6(1):5, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682748.

Green:1954:RAC

[GBF^{+}54] Howard Whipple Green, C. E. Batschelet, W. Thurber Fales, Philip M. Hauser, Lyman L. Hill, Stanley W. Kadow, Earl Lomon Koos, O. E. Lemieux, Harlin G. Loomer, Dorothy W. Myers, Conrad Taeuber, and Leon E. Truesdell. Report of the ASA Committee on Census Enumeration Areas for 1953. The American Statistician, 8(2):21, April/May 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686019.

Gershenson:1951:RAS

[Ger51] M. I. Gershenson. Report of the American Statistical Association representative to the Accident Statistics Committee of the American Standards Association 1950. The American Statistician, 5(4):20–21, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682260.

Gerschenkron:1953:RSI

[Ger53] Alexander Gerschenkron. Reliability of Soviet industrial and national income statistics. The American Statistician, 7(2):18–21, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685633.

Greenberg:1956:FSA

[GGL^{+}56] Leon Greenberg, J. Richard Grant, John W. Lehman, Stanley J. Sigel, and H. E. Riley. Federal statistical activities. The American Statistician, 10(1):4–12, February 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681466.

Gilbert:1958:OIT

[Gil58] E. N. Gilbert. An outline of information theory. The American Statistician, 12(1):13–19, February 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682669.

Gumbel:1954:SAE

[GL54] Emil J. Gumbel and Julius Lieblein. Some applications of extremevalue methods. The American Statistician, 8(5):14–17, December 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681546.

Glaser:1951:IER

[Gla51] Ezra Glaser. Interindustry economics research. The American Statistician, 5(2):9–11, April/May 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681634.

Glaser:1952:RPV

[Gla52] Ezra Glaser. Reservations on Professor Vining’s reflections. The American Statistician, 6(2):27–29, April/May 1952. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681719.

Goldstein:1953:RDS

[Gol53] Harold Goldstein. Research design for studies of the mobility of workers in critical occupations. The American Statistician, 7(1):4– 8, February 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682280.

Greenberg:1959:CCF

[Gre59] B. G. Greenberg. Conduct of cooperative field and clinical trials. The American Statistician, 13(3):13–28, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681365.

Grieves:1958:FSA

[GSP^{+}58] Howard C. Grieves, Peyton Stapp, Jacob Perlman, Robert L. Stein, Rudolph C. Mendelssohn, and Harry Alpert. Federal statistical activities. The American Statistician, 12(3):7–12, June 1958. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682542.

Gulliksen:1952:TMS

[Gul52] Harold Gulliksen. Training in mathematics and statistics desirable for employees at various levels in an educational research organization. The American Statistician, 6(5):8–9, December 1952. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681753.

Haemer:1950:CCP

[Hae50a] Kenneth W. Haemer. Color in chart presentation. The American Statistician, 4(2):20, April/May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681409.

Haemer:1950:SRC

[Hae50b] Kenneth W. Haemer. A simplified ranking chart. The American Statistician, 4(5):21, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681451.

Haemer:1951:PTD

[Hae51] Kenneth W. Haemer. The pseudo third dimension. The American Statistician, 5(4):28, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682265.

Hansen:1952:SSCa

[Han52a] Morris H. Hansen. Statistical standards and the census. The American Statistician, 6(1):7–10, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682751.

Hansen:1952:SSCb

[Han52b] Morris H. Hansen. [statistical standards and the census]: Discussion. The American Statistician, 6(1):13, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682753.

Harrison:1950:JMG

[Har50] Shelby M. Harrison. John M. Glenn. The American Statistician, 4(5):20, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681450.

Hartkemeier:1951:NUD

[Har51] Harry P. Hartkemeier. New undergraduate Degree in Statistics at the University of Missouri. The American Statistician, 5(3):8–9, June/July 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685985.

Hartkemeier:1953:BCH

[Har53] Harry P. Hartkemeier. Binary coding, horizontal card fields and the electronic statistical machine. The American Statistician, 7(4):19– 21, October 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685765.

Harshbarger:1958:SBI

[Har58] Boyd Harshbarger. A survey of both industry and colleges. The American Statistician, 12(1):23–24, February 1958. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682672.

Hayes:1950:SPI

[Hay50] Samuel P. Hayes, Jr. Statistics in the Point IV Program. The American Statistician, 4(1):4–5, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682552.

Hollander:1953:FSA

[HGG^{+}53] Edward D. Hollander, Howard C. Grieves, Charles Gant, W. Parker Mauldin, William Lerner, Helen H. Ringe, and Ida C. Merriam. Federal statistical activities. The American Statistician, 7(2):5–7, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685629.

Hoadley:1951:BES

[Hoa51] Walter E. Hoadley, Jr. The business and economic statistics section. The American Statistician, 5(5):1–13, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681857.

Hoadley:1952:BS

[Hoa52] Walter E. Hoadley, Jr. Benedict Saurino, 1904–1952. The American Statistician, 6(3):33, June/July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685979.

Hoadley:1958:PCa

[Hoa58a] Walter E. Hoadley, Jr. The President’s column. The American Statistician, 12(2):1–29, April 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681507.

Hoadley:1958:PCb

[Hoa58b] Walter E. Hoadley, Jr. The President’s column. The American Statistician, 12(4):1–35, October 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681782.

Hoadley:1958:PCc

[Hoa58c] Walter E. Hoadley, Jr. The President’s column. The American Statistician, 12(5):1–22, December 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682268.

Holubnychy:1958:OSO

[Hol58] Vsevolod Holubnychy. Organization of statistical observation in the U.S.S.R. The American Statistician, 12(3):13–17, June 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682543.

Hotelling:1951:AW

[Hot51] Harold Hotelling. Abraham Wald. The American Statistician, 5 (1):18–19, February 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685920.

Hotelling:1958:SMP

[Hot58] Harold Hotelling. The statistical method and the philosophy of science. The American Statistician, 12(5):9–14, December 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682272.

Huff:1951:LE

[HS51] Darrell Huff and Irving H. Siegel. Letters to the Editor. The American Statistician, 5(5):12–13, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681864.

Hughes:1955:GGS

[Hug55] Everett Cherrington Hughes. The Gleichschaltung of the German Statistical Yearbook. The American Statistician, 9(5):8–11, December 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685502.

Horton:1952:FSA

[HZR^{+}52] Donald C. Horton, Robert Zalkin, Walter F. Ryan, William Lerner, Morris H. Hansen, James W. Turbitt, and Anne Geddes. Federal statistical activities. The American Statistician, 6(3):3–4, June/ July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685972.

Jaffe:1954:OCV

[Jaf54] A. J. Jaffe. One Consumer’s view of the current population survey. The American Statistician, 8(1):11–13, February 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681659.

Johnson:1953:DSP

[Joh53] Palmer O. Johnson. Developments of statistical programs in mid-American universities and colleges. The American Statistician, 7(4):7–15, October 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685763.

Jones:1954:RRC

[Jon54] Homer Jones. Recent revisions of consumer credit statistics. The American Statistician, 8(3):9–35, June/July 1954. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685456.

Jones:1953:FSA

[JWLW53] Homer Jones, Irwin Wolkstein, Emanuel Landau, and Seymour L. Wolfbein. Federal statistical activities. The American Statistician, 7(1):3, February 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682279.

K:1955:SWS

[K.55] F. K. Samuel Weiss, statistician. The American Statistician, 9(4): 12, October 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682115.

Kappel:1951:CSP

[KDOR51] Joseph W. Kappel, Murray D. Dessel, James C. O’Brien, and Ruth Rosenwald. Current statistical projects of the U.S. Government. The American Statistician, 5(1):5, February 1951. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685916.

Kellogg:1953:RCF

[Kel53] Lester S. Kellogg. Report of the Committee on Fellows. The American Statistician, 7(1):1–13, February 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682277.

Kendall:1959:HDE

[Ken59] Maurice G. Kendall. Hiawatha designs an experiment. The American Statistician, 13(5):23–24, December 1959. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682432.

Kish:1959:RPR

[KH59] Leslie Kish and Irene Hess. A “replacement” procedure for reducing the bias of nonresponse. The American Statistician, 13(4):17–19, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685753.

King:1952:AFC

[Kin52] Edmund R. King. Analysis and forecasts of company sales. The American Statistician, 6(3):5–11, June/July 1952. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685973.

Kiser:1952:DDL

[Kis52] Clyde V. Kiser. Discussion of Dr. Linder’s paper. The American Statistician, 6(2):21–22, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681717.

Kimball:1954:LE

[KKP^{+}54] Bradford F. Kimball, Willford I. King, D. E. Paul, Stuart A. Courtis, and Frank A. Hanna. Letters to the Editor. The American Statistician, 8(3):1–18, June/July 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685452.

Klein:1959:NST

[Kle59] Sidney Klein. A note on statistical techniques in Communist China. The American Statistician, 13(3):18–21, June 1959. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681366.

Knapp:1957:NT

[Kna57] Robert H. Knapp. A note on the t test. The American Statistician, 11(1):17, February 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor. org/stable/2681852. See comments [Dwy58].

Koyanagi:1951:SQC

[Koy51] S. Koyanagi. Statistical quality control in Japan. The American Statistician, 5(5):8–9, December 1951. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681861.

Kalabukhov:1958:DLV

[KS58] N. I. Kalabukhov and W. D. C. Scott. A denial of the laws of variation of organisms. The American Statistician, 12(2):20–26, April 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681512.

Lathrop:1954:ORS

[Lat54] John B. Lathrop. Operations research as a science. The American Statistician, 8(3):6–35, June/July 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685455.

Linder:1957:FSA

[LBW^{+}57] Forrest E. Linder, Raymond T. Bowman, Kenneth B. Williams, Jacob S. Siegel, and H. M. Douty. Federal statistical activities. The American Statistician, 11(3):6–12, June 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685678.

Lehman:1954:FSA

[LDL^{+}54] John Lehman, Helen F. Demond, William Lerner, Louise C. Griffith, David L. Kaplan, Abner Hurwitz, Daniel Darling, and H. E. Riley. Federal statistical activities. The American Statistician, 8(4):5–11, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682051.

Leonard:1952:SAU

[Leo52] William R. Leonard. Statistical activities of the United Nations. The American Statistician, 6(2):14, April/May 1952. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681712.

Leonard:1954:PCO

[Leo54] William R. Leonard. [problems of co-ordination in the Canadian Statistical System]: Comments. The American Statistician, 8(4):21–22, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682054.

Lehman:1955:FSAa

[LGD^{+}55] John W. Lehman, J. Richard Grant, Helen F. Demond, Howard C. Grieves, and Vito Natrella. Federal statistical activities. The American Statistician, 9(3):4–5, June 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681456.

Lehman:1955:FSAb

[LGLH55] John Lehman, Paul C. Glick, Louis Levine, and Ethel D. Hoover. Federal statistical activities. The American Statistician, 9(4):6–11, October 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682114.

Likert:1959:PCa

[Lik59a] Rensis Likert. The President’s column. The American Statistician, 13(1):1–30, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685856.

Likert:1959:PCb

[Lik59b] Rensis Likert. The President’s column. The American Statistician, 13(2):1–21, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682307.

Likert:1959:PCc

[Lik59c] Rensis Likert. The President’s column. The American Statistician, 13(3):1–34, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681360.

Likert:1959:PCd

[Lik59d] Rensis Likert. The President’s column. The American Statistician, 13(4):1–46, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685747.

Likert:1959:PCe

[Lik59e] Rensis Likert. The President’s column. The American Statistician, 13(5):1–42, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682425.

Linder:1952:ADS

[Lin52] Forrest E. Linder. Activities of the Demographic and Social Statistics Branch of the Statistical Office of the United Nations. The American Statistician, 6(2):19–21, April/May 1952. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681716.

Littauer:1950:DSQ

[Lit50] S. B. Littauer. The development of statistical quality control in the United States. The American Statistician, 4(5):14–20, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681449.

Littauer:1951:DSQ

[Lit51] S. B. Littauer. The development of statistical quality control in the United States. The American Statistician, 5(1):14–17, February 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685919.

Loftus:1952:IAE

[Lof52] Patrick J. Loftus. International aspects of economic statistics. The American Statistician, 6(2):15–16, April/May 1952. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681713.

Long:1952:SSCa

[Lon52a] Clarence D. Long. [statistical standards and the census]: Discussion. The American Statistician, 6(1):10–13, February 1952. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682752.

Long:1952:SSCb

[Lon52b] Clarence D. Long. [statistical standards and the census]: Rejoinder. The American Statistician, 6(1):13–14, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682754.

Lorimer:1953:NSP

[Lor53] Frank Lorimer. The nature of Soviet population and vital statistics. The American Statistician, 7(2):13–18, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685632.

Lebergott:1957:FSA

[LRP^{+}57] Stanley Lebergott, Walter F. Ryan, John R. Provan, Arnold E. Chase, H. M. Douty, and William Lerner. Federal statistical activities. The American Statistician, 11(4):6–24, October 1957. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681584.

Lehman:1956:FSA

[LYG^{+}56] John W. Lehman, Dudley Young, Edwin D. Goldfield, James B. Eckert, and Nelson M. Bortz. Federal statistical activities. The American Statistician, 10(3):5–25, June 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681772.

Macmillan:1959:RSS

[Mac59] Harold Macmillan. The Royal Statistical Society speech delivered at the 125th Anniversary Dinner. The American Statistician, 13(3):24–25, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681368.

Mahalanobis:1950:SAP

[Mah50] P. C. Mahalanobis. Syllabus for an Advanced Professional Course in Statistical Sampling. The American Statistician, 4(2):13–14, April/ May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681404.

Maher:1958:CAT

[Mah58] John E. Maher. Correlations among time series and Moore’s statistical indicators. The American Statistician, 12(4):15–18, October 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681785.

Mantel:1951:RES

[Man51] Nathan Mantel. Rapid estimation of standard errors of means for small samples. The American Statistician, 5(4):26–27, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682264.

Marshall:1954:PCO

[Mar54] Herbert Marshall. Problems of co-ordination in the Canadian Statistical System. The American Statistician, 8(4):16–21, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682053.

Master:1951:NBP

[MDM51] Arthur M. Master, Louis I. Dublin, and Herbert H. Marks. The normal blood pressure and its clinical implications. The American Statistician, 5(3):6–7, June/July 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685984.

Manvel:1957:FSA

[MDY^{+}57] Allen D. Manvel, Helen F. Demond, Dudley E. Young, Conrad Taeuber, Dana M. Barbour, Forrest E. Linder, Wayne F. Daugherty, and J. Richard Grant. Federal statistical activities. The American Statistician, 11(5):7–12, December 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682564.

Moriyama:1951:CSP

[MGCK51] I. M. Moriyama, Thomas G. Gentel, Herbert S. Conrad, and Morton Kramer. Current statistical projects of the Federal Government. The American Statistician, 5(3):16–17, June/July 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685988.

Mills:1952:RAN

[Mil52] Frederick C. Mills. Report on the activities of the National Bureau of Economic Research. The American Statistician, 6(5):16–17, December 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681759.

Morton:1954:FSA

[MKB^{+}54] J. E. Morton, V. S. Kolesnikoff, Robert W. Burgess, Dudley E. Young, Herman H. Fasteau, and Toshi Toki. Federal statistical activities. The American Statistician, 8(5):3–6, December 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681544.

Martin:1954:FSA

[MLRB54] Margaret E. Martin, John Lehman, Jack B. Robertson, and Howard G. Brunsman. Federal statistical activities. The American Statistician, 8(3):4–5, June/July 1954. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685454.

Mosteller:1950:QAb

[MM50] Frederick Mosteller and Eli S. Marks. Questions and answers. The American Statistician, 4(2):12–15, April/May 1950. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681403.

McKellar:1959:GCSb

[MM59] N. L. McKellar and Isabel McWhinney. Government of Canada statistical activities. The American Statistician, 13(4):13–14, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685751.

Mosteller:1951:QA

[MMS51] Frederick Mosteller, Philip J. McCarthy, and Frederick F. Stephan. Questions and answers. The American Statistician, 5(1):20–21, February 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685922.

Moore:1950:NBE

[Moo50] Geoffrey H. Moore. The National Bureau of Economic Research. The American Statistician, 4(3):13–19, June/July 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682339.

Moore:1954:ABC

[Moo54] Geoffrey H. Moore. Analyzing business cycles. The American Statistician, 8(2):13–19, April/May 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686017.

Moore:1955:DIC

[Moo55] Geoffrey H. Moore. Diffusion indexes: A comment. The American Statistician, 9(4):13–17+30, October 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682116.

Moore:1957:CMD

[Moo57] Geoffrey H. Moore. Comment on “Must the Diffusion Index Lead?”. The American Statistician, 11(4):16–17, October 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681586. See [Val57a].

Morrisey:1950:AMS

[Mor50a] Linda Morrisey. Annual meeting social activities: Especially for wives. The American Statistician, 4(5):9–10, October 1950. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681446.

Morton:1950:STS

[Mor50b] J. E. Morton. Section on the training of statisticians. The American Statistician, 4(2):1, April/May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681399.

Morton:1952:SSD

[Mor52] J. E. Morton. [statistical standards]: Discussion. The American Statistician, 6(1):6–7, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682750.

Mosimann:1950:ITC

[Mos50a] Thomas F. Mosimann. The International Technical Cooperation Program of the Bureau of Labor Statistics — 1945–50. The American Statistician, 4(4):22–23, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685550.

Mosteller:1950:QAc

[Mos50b] Frederick Mosteller. Questions and answers. The American Statistician, 4(4):27–28, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685553.

Mosteller:1950:QAd

[Mos50c] Frederick Mosteller. Questions and answers. The American Statistician, 4(5):7, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681444.

Moss:1951:CSP

[MP51] Milton Moss and Robert B. Pearl. Current statistical projects of the Federal Government. The American Statistician, 5(5):4, December 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681860.

Mosteller:1950:QAa

[MRH50] Frederick Mosteller, P. J. Rulon, and T. E. Harris. Questions and answers. The American Statistician, 4(1):14–15, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682559.

McKellar:1959:GCSa

[MS59] N. L. McKellar and J. A. Sawyer. Government of Canada statistical activities. The American Statistician, 13(3):12–25, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681364.

Mosimann:1959:FSA

[MSBH59] Thomas F. Mosimann, Peyton Stapp, Daniel H. Brill, and Wells Harrington. Federal statistical activities. The American Statistician, 13(5):9–12, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682428.

Martin:1957:FSA

[MSLC57] Margaret E. Martin, Henry D. Sheldon, Stanley Lebergott, and Donald E. Church. Federal statistical activities. The American Statistician, 11(2):6–9, April 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681763.

Murphy:1956:MSU

[Mur56] Walter J. Murphy. Math and statistics and us. The American Statistician, 10(4):17, October 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681325.

Mosteller:1954:RAR

[MWS54] Frederick Mosteller, S. S. Wilks, and Mortimer Spiegelman. Report of ASA representatives to the Social Science Research Council. The American Statistician, 8(5):18, December 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681547.

Martin:1953:FSA

[MYG^{+}53] Margaret E. Martin, Dudley E. Young, Edwin D. Goldfield, Harvey Kailin, Robert J. Myers, James W. McNally, and Dudley E. Young. Federal statistical activities. The American Statistician, 7(5):4–6, December 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685714.

Natrella:1959:NCC

[Nat59] Mary G. Natrella. A note on the computation of χ^{2}. The American Statistician, 13(1):20–21, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685862.

Netschert:1952:MN

[Net52] William Netschert. More news. The American Statistician, 6 (2):9–12, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681708.

Netschert:1955:LE

[Net55] William Netschert. Letter to the Editor. The American Statistician, 9(3):1, June 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681454.

Nicholson:1954:TS

[Nic54] George E. Nicholson, Jr. The training of statisticians. The American Statistician, 8(4):28–32, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682057.

Osborne:1950:CSP

[OCMH50] Harlow A. Osborne, Herbert S. Conrad, Thomas J. Mills, and Harry Hoffner. Current statistical projects of the U.S. Government. The American Statistician, 4(5):4, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681442.

Ostle:1952:QA

[OCSP52] Bernard Ostle, W. S. C., I. R. S., and F. Proschan. Questions and answers. The American Statistician, 6(3):32–33, June/July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685978.

Olds:1957:SSC

[Old57] Edward B. Olds. Social statisticians and the census. The American Statistician, 11(5):26, December 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682567.

Olmstead:1951:STE

[Olm51] Paul S. Olmstead. Statistical training for engineers. The American Statistician, 5(2):15–16, April/May 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681638.

Peterson:1950:LE

[Pet50] Arthur G. Peterson. Letter to the Editor. The American Statistician, 4(1):1, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682550.

Polaneczky:1959:RBO

[Pol59] Aloysius Joseph Polaneczky. Random balance or once upon a time —. The American Statistician, 13(3):22–23, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681367.

Quackenbush:1951:OES

[Qua51] Amanda Quackenbush. Oldest Experiment Station celebrates 75th anniversary. The American Statistician, 5(2):11–12, April/May 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681635.

Rothman:1959:CMA

[RB59] Abe Rothman and Edgar M. Bisgyer. Characteristics of the membership of the American Statistical Association, 1958. The American Statistician, 13(4):31–37, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685755.

Rubin:1957:QAe

[RD57] Ernest Rubin and Philip Desind. Questions and answers. The American Statistician, 11(5):23–25, December 1957. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682566.

Rubin:1955:QAa

[RE55] Ernest Rubin and Libert Ehrman. Questions and answers. The American Statistician, 9(1):24–25, February 1955. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682353.

Rubin:1958:QAa

[RH58] Ernest Rubin and Q. M. Hussain. Questions and answers. The American Statistician, 12(1):28–30, February 1958. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682674.

Rubin:1959:QAd

[RH59a] Ernest Rubin and Britton Harris. Questions and answers. The American Statistician, 13(4):38–46, October 1959. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685756.

Rubin:1959:QAc

[RH59b] Ernest Rubin and Robert G. Hoffman. Questions and answers. The American Statistician, 13(3):27–28, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681369.

Rubin:1959:QAe

[RH59c] Ernest Rubin and Frank Holzman. Questions and answers. The American Statistician, 13(5):19–27, December 1959. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682431.

Rice:1950:NLF

[Ric50] Stuart A. Rice. New legislation on federal statistical coordination. The American Statistician, 4(5):8–9, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681445.

Rice:1951:RBP

[Ric51] Stuart A. Rice. Revision of base period for general-purpose indexes. The American Statistician, 5(4):18–19, October 1951. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682258.

Rice:1952:SCS

[Ric52] Stuart A. Rice. Statistical conceptions in the Soviet Union examined from generally accepted scientific viewpoints. The American Statistician, 6(3):12–15, June/July 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685974.

Rice:1953:I

[Ric53a] Stuart A. Rice. Introduction. The American Statistician, 7(2):8, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685630.

Rice:1953:SSF

[Ric53b] Stuart A. Rice. Statistics and social forces. The American Statistician, 7(5):7–11, December 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685715.

Rice:1959:WFO

[Ric59] Stuart A. Rice. William Fielding Ogburn, 1886–1959. The American Statistician, 13(3):29, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681370.

Riffenburgh:1959:NGM

[Rif59] Robert H. Riffenburgh. A note on the graphic method for obtaining a least squares fit. The American Statistician, 13(1):21, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685863.

Rubin:1955:QAb

[RJ55] Ernest Rubin and Naum Jasny. Questions and answers. The American Statistician, 9(2):22–24, April 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681972.

Rubin:1957:QAa

[RJ57a] Ernest Rubin and Naum Jasny. Questions and answers. The American Statistician, 11(1):18–20, February 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681853.

Rubin:1957:QAd

[RJ57b] Ernest Rubin and Naum Jasny. Questions and answers. The American Statistician, 11(4):18–20, October 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681588.

Rose:1958:FSA

[RJKP58] J. Hugh Rose, Sidney A. Jaffe, Joseph W. Kappel, and Rexford C. Parmelee. Federal statistical activities. The American Statistician, 12(4):9–14, October 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681784.

Riley:1955:FSA

[RLV^{+}55] Donald C. Riley, John Lehman, Virginia T. Venneman, Donald E. Church, Carl A. S. Coan, and W. Duane Evans. Federal statistical activities. The American Statistician, 9(2):7–9, April 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681968.

Rosander:1950:NHS

[Ros50] A. C. Rosander. A note on high school courses in quantitative thinking. The American Statistician, 4(3):16–17, June/July 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682340.

Rottenberg:1958:TSC

[Rot58] Simon Rottenberg. The teaching of statistics in Chile. The American Statistician, 12(2):14–15, April 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681510.

Rubin:1955:QAd

[RR55] Ernest Rubin and Morton S. Raff. Questions and answers. The American Statistician, 9(5):19–21, December 1955. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685504.

Rubin:1955:QAc

[RS55] Ernest Rubin and John H. Smith. Questions and answers. The American Statistician, 9(3):21–22, June 1955. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681460.

Rubin:1956:QAc

[RS56] Ernest Rubin and John H. Smith. Questions and answers. The American Statistician, 10(3):23–24, June 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681776.

Rubin:1958:QAd

[RS58] Ernest Rubin and John H. Smith. Questions and answers. The American Statistician, 12(4):24–41, October 1958. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681787.

Rubin:1959:QAa

[RSPB59] Ernest Rubin, Galina V. Selegen, Victor P. Petrov, and James W. Brackett. Questions and answers. The American Statistician, 13 (1):14–19, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685860.

Ringe:1952:CSP

[RSS^{+}52] Helen H. Ringe, Allan D. Searle, Walt Simmons, S. Shapiro, Boyd Ladd, H. M. Douty, and Bernard B. Watson. Current statistical projects of the Federal Government. The American Statistician, 6(4):4–27, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682695.

Rubin:1954:QAa

[Rub54a] Ernest Rubin. Questions and answers. The American Statistician, 8(3):34–35, June/July 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685461.

Rubin:1954:QAb

[Rub54b] Ernest Rubin. Questions and answers. The American Statistician, 8(4):33–34, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682058.

Rubin:1954:QAc

[Rub54c] Ernest Rubin. Questions and answers. The American Statistician, 8(5):19–21, December 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681548.

Rubin:1956:QAa

[Rub56a] Ernest Rubin. Questions and answers. The American Statistician, 10(1):14–16, February 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681470.

Rubin:1956:QAb

[Rub56b] Ernest Rubin. Questions and answers. The American Statistician, 10(2):15–16, April 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681600.

Rubin:1956:QAd

[Rub56c] Ernest Rubin. Questions and answers. The American Statistician, 10(4):18–23, October 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681326.

Rubin:1956:QAe

[Rub56d] Ernest Rubin. Questions and answers. The American Statistician, 10(5):20–21, December 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681700.

Rubin:1957:QAb

[Rub57a] Ernest Rubin. Questions and answers. The American Statistician, 11(2):22, April 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681767.

Rubin:1957:QAc

[Rub57b] Ernest Rubin. Questions and answers. The American Statistician, 11(3):31–32, June 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685681.

Rubin:1958:QAb

[Rub58a] Ernest Rubin. Questions and answers. The American Statistician, 12(2):21–22, April 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681513.

Rubin:1958:QAc

[Rub58b] Ernest Rubin. Questions and answers. The American Statistician, 12(3):21–24, June 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682545.

Rubin:1958:QAe

[Rub58c] Ernest Rubin. Questions and answers. The American Statistician, 12(5):16–17, December 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682273.

Rubin:1959:QAb

[Rub59] Ernest Rubin. Questions and answers. The American Statistician, 13(2):23–24, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682314.

Sagen:1953:PVS

[Sag53] O. K. Sagen. Production of vital statistics as a combined Federal– State operation. The American Statistician, 7(4):16–18, October 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685764.

Sarhan:1957:TSE

[Sar57] A. E. Sarhan. Teaching of statistics in Egypt. The American Statistician, 11(1):15–17, February 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681851.

Schwartz:1953:OOS

[Sch53] Harry Schwartz. The organization and operation of the Soviet statistical apparatus. The American Statistician, 7(2):9–13, April/ May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685631.

Sawyer:1959:GCS

[SCJ59] J. A. Sawyer, E. B. Carty, and D. H. Jones. Government of Canada statistical activities. The American Statistician, 13(5):13, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682429.

Stauber:1959:FSA

[SDF^{+}59] B. Ralph Stauber, H. M. Douty, Willard Fazar, Richard H. Jordan, William Weinfeld, and Allen D. Manvel. Federal statistical activities. The American Statistician, 13(2):9–12, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682310.

Shiskin:1955:SCU

[Shi55] Julius Shiskin. Seasonal computations on Univac. The American Statistician, 9(1):19–23, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682352.

Sholes:1959:AMC

[Sho59] De Ver Sholes. Annual Midwest Conference in Chicago. The American Statistician, 13(1):22–23, February 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685864.

Siegel:1950:RPR

[Sie50] Irving H. Siegel. Report on plans for regional meetings. The American Statistician, 4(1):18, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682561.

Siegel:1952:LE

[Sie52] Irving H. Siegel. Letters to the Editor. The American Statistician, 6(1):1–21, February 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682746.

Siegel:1957:NS

[Sie57] Sidney Siegel. Nonparametric statistics. The American Statistician, 11(3):13–19, June 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685679.

Simmons:1956:MIC

[Sim56] Walt R. Simmons. Measurement of industrial and commercial employment in the United States. The American Statistician, 10 (4):11–16, October 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681324.

Spurr:1951:L

[SK51] William A. Spurr and John W. Kendrick. Letters. The American Statistician, 5(4):21, October 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682261.

Shiskin:1958:FSA

[SK58] Julius Shiskin and Joseph W. Kappel. Federal statistical activities. The American Statistician, 12(5):7–8, December 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682271.

Sigel:1959:FSA

[SLCM59] Stanley J. Sigel, Forrest E. Linder, Vladimir D. Chavrid, and Margaret E. Martin. Federal statistical activities. The American Statistician, 13(3):6–11, June 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681363.

Sirken:1958:FSA

[SLK^{+}58] Monroe G. Sirken, Forrest E. Linder, Harry W. Ketchum, John Storck, William Lerner, Lelia M. Easson, Louis Levine, and Conrad Taeuber. Federal statistical activities. The American Statistician, 12(2):8–13, April 1958. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681509.

Steffes:1956:FSA

[SLW^{+}56] Robert B. Steffes, Louis Levine, Seymour Wolfbein, George St. J. Perrott, and Kenneth B. Williams. Federal statistical activities. The American Statistician, 10(4):5–10, October 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681323.

Smalley:1954:EDO

[Sma54] C. W. Smalley. Estimating daily order receipts from weight of mail. The American Statistician, 8(1):14–15, February 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681660.

Smith:1956:LE

[Smi56] H. Fairfield Smith. Letter to the Editor. The American Statistician, 10(4):1, October 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681321.

Schwartz:1954:FSA

[SMW^{+}54] M. H. Schwartz, Allen D. Manvel, Irwin Wolkstein, William Lerner, and Harvey Kailin. Federal statistical activities. The American Statistician, 8(1):4–5, February 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681657.

Steinberg:1954:FSA

[SPE54] Joseph Steinberg, Rexford C. Parmelee, and Grover Ensley. Federal statistical activities. The American Statistician, 8(2):5–6, April/ May 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686015.

Sprowls:1956:ACF

[Spr56] R. Clay Sprowls. An automatic coin flipper for sampling demonstrations. The American Statistician, 10(2):12–13, April 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681598.

Sprowls:1959:AST

[Spr59] R. Clay Sprowls. An answer sheet for teaching tests of hypotheses. The American Statistician, 13(2):18, April 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682312.

Spurr:1952:RCS

[Spu52] William A. Spurr. Recent changes in statistics courses and curricula in Western universities. The American Statistician, 6(4):15–17, October 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682699.

Nations:1950:PSS

[Sta50] Statistical Office of the United Nations. The preparation of sampling survey reports. The American Statistician, 4(3):6–12, June/ July 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682337.

Stapleton:1959:LE

[Sta59] J. H. Stapleton. Letter to the Editor. The American Statistician, 13(4):45, October 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685759.

Stephan:1955:SQM

[Ste55a] Frederick F. Stephan. Some questions on the measurement of unemployment. The American Statistician, 9(1):8–9, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682348.

Stewart:1955:USE

[Ste55b] Charles D. Stewart. Unemployment statistics and economic policy uses. The American Statistician, 9(1):10–14, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682349.

Sturtevant:1950:SCP

[Stu50] John V. Sturtevant. The scatter chart as a practical tool. A simple means of presenting joint two-way influence. The American Statistician, 4(3):20–21, June/July 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682342.

Stapp:1953:FSA

[SWHS53] Peyton Stapp, Helen Wood, Thomas G. Hutton, and Lawrence A. Sternberg. Federal statistical activities. The American Statistician, 7(4):3–6, October 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685762.

Steinberg:1956:FSA

[SWL56] Joseph Steinberg, Seymour Wolfbein, and Louis Levine. Federal statistical activities. The American Statistician, 10(2):5–11, April 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681596.

Spiegelman:1953:RAR

[SWW53] Mortimer Spiegelman, S. S. Wilks, and Holbrook Working. Report of ASA representatives to the Social Science Research Council. The American Statistician, 7(2):22, April/May 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685634.

Tepping:1952:DUT

[Tep52] Benjamin J. Tepping. A desirable undergraduate training program in statistics. The American Statistician, 6(5):9–11, December 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681754.

Teper:1955:USE

[Tep55] Lazare Teper. [unemployment statistics and economic policy uses]: Discussion. The American Statistician, 9(1):15–17, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682351.

Thomas:1953:EG

[Tho53] Woodlief Thomas. Emanuel A. Goldenweiser. The American Statistician, 7(3):6–7, June/July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685492.

Ullman:1959:CUR

[Ull59] Morris B. Ullman. Content of the 1959 U.S.S.R. census of population. The American Statistician, 13(5):14–24, December 1959. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682430.

Valavanis:1957:MDIa

[Val57a] Stefan Valavanis. Must the diffusion index lead? The American Statistician, 11(4):12–16, October 1957. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681585. See comment [Moo57].

Valavanis:1957:MDIb

[Val57b] Stefan Valavanis. [must the diffusion index lead?]: Rejoinder. The American Statistician, 11(4):17, October 1957. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681587.

Vining:1952:RUR

[Vin52] Rutledge Vining. Reflections upon reading Mr. Ezra Glaser’s description of interindustry economics research. The American Statistician, 6(2):23–26, April/May 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681718.

Volin:1953:ASS

[Vol53] Lazar Volin. Agricultural statistics in Soviet Russia: Their usability and reliability. The American Statistician, 7(3):8–12, June/July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685493.

Walker:1951:SLS

[Wal51a] Helen M. Walker. Statistical literacy in the social sciences. The American Statistician, 5(1):6–12, February 1951. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685917.

Wallis:1951:SC

[Wal51b] W. Allen Wallis. Statistics at Chicago. The American Statistician, 5(2):16–17, April/May 1951. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681639.

Wasserman:1958:TUS

[Was58] William Wasserman. The teaching and use of statistics in Turkey. The American Statistician, 12(2):16–18, April 1958. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681511.

Whipple:1956:RCC

[WBB^{+}56] Howard Whipple, C. E. Batschelet, Howard G. Brunsman, F. A. Buechel, Sidney Cohen, Florence E. Cuttrell, Philip M. Hauser, Stanley W. Kadow, O. A. Lemieux, Harlin G. Loomer, Ferdinand F. Mauser, W. E. McCann, Dorothy W. Myers, Conrad Taeuber, and Leon E. Truesdell. Report of the Committee on Census Enumeration Areas, 1955. The American Statistician, 10(2):16–17, April 1956. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681601.

Walker:1950:MAT

[WD50] Helen M. Walker and Walter N. Durost. A model to aid in teaching partial correlation. The American Statistician, 4(5):5–7, October 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681443.

Weiss:1950:RS

[Wei50] Samuel Weiss. The role of the statistician. The American Statistician, 4(3):1, June/July 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682335.

Weiss:1954:PAL

[Wei54] Samuel Weiss. Proposed amendment to the by-laws on election of Fellows. The American Statistician, 8(1):1, February 1954. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681655.

Wickens:1954:PFS

[Wic54] Aryness Joy Wickens. Proposals for a Federal Statistical System in the United States. The American Statistician, 8(4):12–16, October 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682052.

Wixon:1952:RFP

[Wix52] Rufus Wixon. Report on Financial Planning and Control Section. The American Statistician, 6(5):13–14, December 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681757.

Woytinsky:1954:LMS

[Woy54] W. S. Woytinsky. Limits of mathematics in statistics. The American Statistician, 8(1):6–18, February 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681658.

Weiss:1955:FMS

[WR55] Samuel Weiss and Abe Rothman. The financial and membership status of the American Statistical Association. The American Statistician, 9(2):10–14, April 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681969.

Wright:1952:ICS

[Wri52] C. Ashley Wright. Intra-company statistics in the Standard Oil Company (New Jersey). The American Statistician, 6(5):4–7, December 1952. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681752.

Watkins:1954:RCE

[WRW54] Ralph J. Watkins, Donald C. Riley, and Samuel S. Wilks. Report of the Committee on Elections. The American Statistician, 8(2):1, April/May 1954. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686013.

Working:1950:PSE

[WW50] Holbrook Working and Ralph J. Watkins. Program of the Section on Economic and Business Statistics. The American Statistician, 4(1):9–10, February 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682555.

Yates:1950:SPS

[Yat50] F. Yates. Syllabus for professional statisticians who intend to become experts in sampling census and survey methods. The American Statistician, 4(2):14–15, April/May 1950. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681405.

Young:1955:FSA

[YHY^{+}55] Ralph A. Young, Abner Hurwitz, Dudley E. Young, George Jaszi, H. E. Riley, Walter W. Schneider, Paul Shapiro, and William Lerner. Federal statistical activities. The American Statistician, 9(1):5–7, February 1955. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682347.

Young:1953:FSA

[YLMM53] Dudley E. Young, William Lerner, Allen D. Manvel, and James L. McPherson. Federal statistical activities. The American Statistician, 7(3):4–5, June/July 1953. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685491.